### v3

# Ramification filtration via deformations, II

Abstract.
Let $\mathcal{K}$ be a field of formal Laurent series with coefficients in a finite field of characteristic $p$. For $M\in\mathbb{N}$, let $\mathcal{G}_{<p,M}$ be the maximal quotient of the Galois group of $\mathcal{K}$ of period $p^{M}$ and nilpotent class $<p$ and $\{\mathcal{G}_{<p,M}^{(v)}\}_{v\geqslant 0}$ — the filtration by ramification subgroups in upper numbering. Let $\mathcal{G}_{<p,M}=G(\mathcal{L})$ be the identification of nilpotent Artin-Schreier theory: here $G(\mathcal{L})$ is the group obtained from a suitable profinite Lie $\mathbb{Z}/p^{M}$-algebra $\mathcal{L}$ via the Campbell-Hausdorff composition law. We develop new techniques to obtain a “geometrical” construction of the ideals $\mathcal{L}^{(v)}$ such that $G(\mathcal{L}^{(v)})=\mathcal{G}_{<p,M}^{(v)}$. Given $v_{0}\geqslant 1$, we construct a decreasing central filtration $\mathcal{L}(w)$, $1\leqslant w\leqslant p$, on $\mathcal{L}$, an epimorphism of Lie $\mathbb{Z}/p^{M}$-algebras $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}:=\mathcal{L}/\mathcal{L}(p)$, and a unipotent action $\Omega$ of $\mathbb{Z}/p^{M}$ on $\bar{\mathcal{L}}^{{\dagger}}$, which induces the identity action on $\bar{\mathcal{L}}$. Suppose $d\Omega=B^{{\dagger}}$, where $B^{{\dagger}}\in\operatorname{Diff}\bar{\mathcal{L}}^{{\dagger}}$, and ¯$\bar{\mathcal{L}}^{{\dagger}[v_{0}]}$ is the ideal of $\bar{\mathcal{L}}^{{\dagger}}$ generated by the elements of $B^{{\dagger}}(\bar{\mathcal{L}}^{{\dagger}})$. Our main result states that the ramification ideal $\mathcal{L}^{(v_{0})}$ appears as the preimage of the ideal in $\bar{\mathcal{L}}$ generated by $\bar{\mathcal{V}}B^{{\dagger}}(\bar{\mathcal{L}}^{{\dagger}[v_{0}]})$. In the last section we apply this to the explicit construction of generators of ¯$\bar{\mathcal{L}}^{(v_{0})}$. The paper justifies a geometrical origin of ramification subgroups of $\Gamma_{K}$ and can be used for further developing of non-abelian local class field theory.
Key words and phrases:
local field, Galois group, ramification filtration
2010 Mathematics Subject Classification:
11S15, 11S20
Introduction
Everywhere in the paper $p$ is a prime number. If $K$ is a field we denote by $K_{sep}$ its separable closure and set $\Gamma_{K}=\mathrm{Gal}(K_{sep}/K)$.
0.1. Motivation
We say that $K$ is cdvf if it is a complete discrete valuation field with finite residue field of characteristic $p$. Then the structure of the group $\Gamma_{K}$ , cf. [18, 29], can be explicitly described. As a matter of fact, it depends only on $\mathrm{char}K$ and the torsion subgroup of $K^{*}$. In particular, it is very weak invariant of $K$. Note that in the case of algebraic number fields $E$ the situation is quite opposite. Here by the Neukirch-Uchida theorem, the functor $E\mapsto\mathrm{Gal}(E_{sep}/E)$ is fully faithful [27]. In other words, $\Gamma_{E}$ is an absolute invariant of the field $E$.
For a cdvf $K$, consider the decreasing filtration by normal ramification subgroups $\{\Gamma_{K}^{(v)}\}_{v\geqslant 0}$ in $\Gamma_{K}$, [28]. This (ramification) filtration allows us to recover arithmetic invariants (say, discriminants, differentes, etc) of subextensions $L$ of $K$ in $K_{sep}$. The importance of ramification filtration was never doubted because this is quite non-trivial number theoretic interpretation of the fundamental geometric concept of ramification. However, during very long time the structure of ramification subgroups $\Gamma_{K}^{(v)}$ was known just modulo the subgroup of second commutators of $\Gamma_{K}$. (In other words, we knew only the images of $\Gamma_{K}^{(v)}$ in the maximal abelian quotient $\Gamma_{K}^{ab}$ of $\Gamma_{K}$ given by local class field theory.) Only in [26] Mochizuki proved that in the mixed characteristic case $\Gamma_{K}$ together with additional structure given by its ramification filtration behaves like a Galois group of algebraic number field. In other words, if $K$ is a cdvf of characteristic 0 then the functor $K\mapsto(\Gamma_{K},\{\Gamma_{K}^{(v)}\}_{v\geqslant 0})$ is fully faithful. This gave the first formal confirmation of importance of ramification theory in the number theoretic setting. As a result, the ramification filtration should not anymore treated just as a technical concept, it is strict invariant of $K$ and, as such, requires careful and detailed study. Notice that Mochizuki’s proof is a spectacular application of the theory of $p$-adic representations; at the same time his proof is highly non-efficient.
It is well-known that the group structure of $\Gamma_{K}$ can be easily recovered from its maximal pro-$p$-quotient $\Gamma_{K}(p)=\mathrm{Gal}(K(p)/K)$, where $K(p)$ is the maximal $p$-extension of $K$ in $K_{sep}$, cf. [18]. This group is provided with induced ramification filtration $\{\Gamma_{K}(p)^{(v)}\}_{v\geqslant 0}$. In [1, 2, 3] the author obtained first explicit description (at substantially non-abelian level) of the subgroups $\Gamma_{K}(p)^{(v)}$ modulo the subgroup $C_{p}(\Gamma_{K}(p))$ of commutators of order $\geqslant p$ in $\Gamma_{K}(p)$. This approach systematically used the equivalence of the categories of $p$-groups and Lie $\mathbb{Z}_{p}$-algebras with nilpotent class $<p$ and new techniques which allowed us to work with ramification properties in terms of the appropriate Lie algebras. These results already found the following applications.
0.1.1. Mixed characteristic case
In [9, 10, 11] we used the Fontaine-Wintenberger field-of-norms functor [33] to extend the approach from [1, 2, 3] to the case of cdvf’s $K$ of mixed characteristic under the assumption that $K$ contains non-trivial $p$-th roots of unity. In particular, suppose that $K$ contains a primitive $p^{M}$-th root of unity, $M\geqslant 1$, and let $K(s,M)$ be the maximal $p$-extension of $K$ in $K(p)$ with the Galois group of period $p^{M}$ and nilpotent class $s<p$. Then we obtained explicit value for the number $v(s,M)$ such that $\mathrm{Gal}(K(s,M)/K)^{(v)}=e$ iff $v>v(s,M)$.
0.1.2. Local analogue of the Grothendieck Conjecture
In [4, 7] we obtained quite efficient generalisation of Mochizuki’s result [26] in the context of $p$-extensions for cdvf $K$ of any characteristic. Namely, if $K$ is cdvf then the functor $K\mapsto(\Gamma_{K}(p),\{\Gamma_{K}(p)^{(v)}\}_{v\geqslant 0})$ is fully faithful. If $p\geqslant 3$ the proof of the equicharacteristic case in [4] uses explicit description of the subgroups $\Gamma_{K}(p)^{(v)}$ modulo third commutators $C_{3}(\Gamma_{K}(p))\supset C_{p}(\Gamma_{K}(p))$. Then the mixed chracteristic case is deduced via application of the field-of-norms functor. In the case $p=2$ our approach does not work because modulo $C_{p}(\Gamma_{K}(p))=C_{2}(\Gamma_{K}(2))$ we are just in the abelian situation where the ramification filtration does not carry essential information. However, in [7] we studied the images of ramification filtration in the Magnus algebra associated with $\Gamma_{K}(2)$ modulo subgroup $C_{3}(\Gamma_{K}(2))$ of third commutators. This new technique allowed us to prove the case $p=2$ in the both (equicharacteristic and mixed characteristic) cases. Our method even worked in the case of higher local fields, cf. [5, 6].
0.1.3. “Small” $p$-adic representations
Suppose $V$ is a finite $\mathbb{Z}_{p}$-module and $\pi:\Gamma_{K}\longrightarrow GL_{\mathbb{Z}_{p}}(V)$ is a group homomorphism. Then we have a finite field extension $K_{\pi}:=K_{sep}^{\mathrm{Ker}\pi}$ of $K$ and arithmetic properties of $K_{\pi}/K$ depend on the images $\pi(\Gamma_{K}^{(v)})=\mathrm{Gal}(K_{\pi}/K)^{(v)}\subset GL_{\mathbb{Z}_{p}}(V)$ of subgroups $\Gamma_{K}^{(v)}$ of $\Gamma_{K}$.
Question. Describe the subgroups $\pi(\Gamma_{K}^{(v)})\subset GL_{\mathbb{Z}_{p}}(V)$.
In [13] we introduced the concept of “small” $p$-adic representation. By definition, $\pi$ is small if for the wild inertia subgroup $\tilde{I}_{K}$ of $\Gamma_{K}$ it holds $\pi(\tilde{I}_{K})=\exp(L)$, where $L$ is a Lie subalgebra in $\mathrm{End}_{\mathbb{Z}_{p}}V$ such that $L^{p}=0$. We proved that if $\mathrm{char}K=p$ and $\pi$ is small then all images $\pi(\Gamma_{K}^{(v)})\subset GL_{\mathbb{Z}_{p}}(V)$ can be uniquely recovered from explicitly constructed differential form on Fontaine’s $\varphi$-module $\mathcal{M}$ associated with $\pi$. The construction of such differential form is purely geometric. It is related to the canonical connection $\nabla$ on $\mathcal{M}$ and its preimages $\varphi^{*N}(\nabla)$ on $\mathcal{M}\otimes_{K}K_{rad}$, where $K_{rad}$ is the radical closure of $K$ in $K_{sep}$. In particular, this construction gives quite unexpected relation of the number-theoretic ramification theory to a very popular theory of $\mathcal{D}$-modules, lifts of Frobenius, etc.
0.1.4. Non-abelian explicit reciprocity laws (ERL)
Suppose that $V$ is a finite cyclic $\mathbb{Z}_{p}$-module. If $V=(\mathbb{Z}/p^{M+1})v^{*}$ then $\mathrm{Gal}(K_{\pi}/K)\subset(\mathbb{Z}/p^{M+1})^{\times}$ and $\pi$ factors as
$\Gamma_{K}\longrightarrow\Gamma_{K}^{ab}\longrightarrow\mathrm{Gal}(K_{\pi}/K)\subset(\mathbb{Z}/p)^{\times}\times\mathbb{Z}/p^{M}\,.$
Since $K_{\pi}/K$ is abelian all $\mathrm{Gal}(K_{\pi}/K)^{(v)}$ come from $\Gamma_{K}^{ab(v)}$ which are well-known by local class field theory. Then the corresponding images of $\Gamma_{K}^{(v)}$ can be found via one-or-another ERL for the Hilbert symbol. For example, suppose $K=k((t_{0}))$ contains a primitive $p^{M}$-th root of unity and $[K_{\pi}:K]=p^{M}$ then $K_{\pi}=K(\root p^{M} \of{v_{\pi}})$, $v_{\pi}\in K^{*}$. So the images $\pi(\Gamma_{K}^{(v)})$ can be calculated via calculations with the $M$-th Hilbert symbols $(u,v_{\pi})$, where $u\equiv 1\,\mathrm{mod}\,t_{0}^{v}$. The study of different ERL’s for the Hilbert symbol (and its analogs) is one of most popular areas of Algebraic Number Theory. It has very rich history [25] started with Kummer’s work in 1859. Known ERL’s have more-or-less similar structure: the answer appears as a residue of suitable explicitly constructed differential form. As example, cf. fundamental contributions of Shafarevich [30], Artin-Hasse [14], Iwasawa [20], Brückner-Vostokov [15, 32], Witt [17], Coates-Wiles [16], et al. If $V$ from 0.1.3 is not cyclic one-or-another non-abelian version of the Hilbert symbol and the corresponding ERL’s could be used to study the question from Section 0.1.3 in full generality. Such symbol should use as essential ingredients nilpotent versions of local class field theory and of Kummer or Artin-Schreier theory. The idea of creating nilpotent class field theory was very popular for a very long time, e.g. [30, 22, 21]. Recently, F. Laubie [23] obtained such theory by generalizing the approach from [22]. It is based on the idea of iteration of Lubin-Tate theories via the field-of-norms functor. This theory posseses many nice properties but is not adjusted for obtaining (even very well-known) explicit information about involved Galois groups. A nilpotent analog of the Artin-Schreier theory was constructed by the author in [1, 2]; it allowed us to obtain new results about ramification filtration in [1, 3]. Introduction of nilpotent analog of Hilbert symbol could be stated now as the question about linking the Laubie and our theories. This will allow us to benefit from functorial properties of the first and explicit constructions of the second theories. Notice that the case of “small” represntations [13] gives a positive feeling about existence of ERL’s in the non-abelian context.
0.1.5. Geometrical approach to ramification filtration
The above applications not only illustrate a fundamental role of explicit results from [1, 2, 3] but also rise the question about their geometric origin and more structural approach. The proofs in [1] are reasonably technical but their modulo $p^{M}$-version in [2, 3] relies on heavy computations with enveloping algebras. When working with Lie algebras of nilpotent class $<p$ the most efficient and clear approach should use one-or-another deformation theory. Such theory was recently constructed in the case of Galois groups of period $p$ in [12] and this gave us geometrical approach to [1]. In this paper we construct more general deformation theory which adjusted to the modulo $p^{M}$ situation. This theory not just gives the results of [2, 3] but also considerably simplifies the approach from [12]. We expect that this new techniques could be applied to develop the modulo $p^{M}$-case of results obtained in the mixed characteristic case in [9, 10]. Note that in [12] we gave also an explicit description of arbitrarily long segments of ramification filtration (not just of a single ramification subgroup) in the case of Galois groups of period $p$. This description admits a straight generalisation to the modulo $p^{M}$-situation. As a result, we have a substantial control on explicit information about arithmetical structure of Galois groups of cdvf’s. We also expect the existence of other fundamental relations between Arithmetic and Geometry of cdvf’s.
0.2. The content
Everywhere in the paper $M\in\mathbb{N}$ and $v_{0}\in\mathbb{R}_{>0}$ are fixed. For $s\in\mathbb{N}$ and a topological group $G$, we denote by $C_{s}(G)$ the closure of the subgroup of $G$ generated by the commutators of order $\geqslant s$. If $L$ is a toplogical Lie algebra then $C_{s}(L)$ is the closure of the ideal generated by commutators of degree $\geqslant s$. For any topological $A$-module $M$ and $B$, we use the notation $M_{B}:=M\hat{\otimes}_{A}B$.
Let $\mathcal{K}$ be a cdvf and let $\mathcal{K}_{<p,M}$ be a maximal $p$-extension of $\mathcal{K}$ with the Galois group $\mathrm{Gal}(\mathcal{K}_{<p,M}/\mathcal{K}):=\mathcal{G}_{<p,M}$ of nilpotence class $<p$ and exponent $p^{M}$. If $L$ is a Lie $\mathbb{Z}/p^{M}$-algebra then $G(L)$ is the group obtained from the elements of $L$ via the Campbell-Hausdorff composition law $\circ$, cf. Sect. 1.2.
Consider a decreasing filtration by ramification subgroups in the upper numbering $\{\mathcal{G}^{(v)}_{<p,M}\}_{v\geqslant 0}$ of $\mathcal{G}_{<p,M}$. In [12] we constructed a deformation theory which allowed us to retrieve (in essentially geometrical way) the results from [1]. In this paper, we obtain a generalisation (and also a considerable simplification) of our approach from [12] to the case of Galois groups $\mathcal{G}_{<p,M}$ with arbitrary exponent $p^{M}$. We also explain how this theory can be applied to recover the structure of the ramification subgroups $\mathcal{G}_{<p,M}^{(v)}$. Our approach could be overviewed as follows.
a) Let $\mathcal{K}=k((t_{0}))$, where $k\simeq\mathbb{F}_{p^{N_{0}}}$. Then we can use the uniformiser $t_{0}$ to define the lifts modulo $p^{M}$, $O_{M}(\mathcal{K})=W_{M}(k)((t))\subset W_{M}(\mathcal{K})$ and $O_{M}(\mathcal{K}_{sep})\subset W_{M}(\mathcal{K}_{sep})$. Here $W_{M}$ is the functor of Witt vectors of length $M$ and $t=[t_{0}]\in W_{M}(\mathcal{K})$ is the Teichmuller representative of $t_{0}$. This allows us to apply the nilpotent version of Artin-Shreier theory, cf. [2], to fix an identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$, where $\mathcal{L}$ is explicitly constructed pro-finite Lie $\mathbb{Z}/p^{M}$-algebra. The identification $\eta^{0}$ appears via choosing a suitable $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$, taking $f\in\mathcal{L}_{O_{M}(\mathcal{K}_{sep})}$ such that $\sigma(f)=e\circ f$ and setting for any $\tau\in\mathcal{G}_{<p,M}$, $\eta^{0}(\tau)=(-f)\circ\tau(f)\in G(\mathcal{L})$. This provides us also with “ramification” ideals $\mathcal{L}^{(v)}$ in $\mathcal{L}$ such that $\eta^{0}(\mathcal{G}_{<p,M}^{(v)})=\mathcal{L}^{(v)}$.
b) The extension of scalars $\mathcal{L}_{W_{M}(k)}$ appears with the set of explicit generators $\{D_{an}\ |\ \mathrm{gcd}(a,p)=1,n\in\mathbb{Z}/N_{0}\}\cup\{D_{0}\}$. We use them to define a decreasing filtration of ideals $\mathcal{L}=\mathcal{L}(1)\supset\mathcal{L}(2)\supset\ldots\supset\mathcal{L}(p)$ and provide ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$ with induced filtration ¯$\bar{\mathcal{L}}(w)$, $1\leqslant w\leqslant p$. The algebra $\bar{\mathcal{L}}$ is considerably smaller than $\mathcal{L}$; it is generated over $W_{M}(k)$ by all $D_{an}$ with $0\leqslant a<(p-1)v_{0}$. However, we prove that $\mathcal{L}(p)\subset\mathcal{L}^{(v_{0})}$ and, as a result, it will be sufficient to describe the image ¯$\bar{\mathcal{L}}^{(v_{0})}$ of $\mathcal{L}^{(v_{0})}$ in $\bar{\mathcal{L}}$.
c) Our next step is a construction of a new Lie $\mathbb{Z}/p^{M}$-algebra $\bar{\mathcal{L}}^{{\dagger}}$ together with projection $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}$. As a matter of fact, there is a splitting of Lie algebras ¯$\bar{\mathcal{L}}^{{\dagger}}=\mathrm{Ker}\bar{\mathcal{V}}\oplus\bar{\mathcal{L}}$. The construction of $\bar{\mathcal{L}}^{{\dagger}}$ depends on a construction of a special system of coefficients inside ${O_{M}(\mathcal{K})}$, cf. Sect. 2. We just mention that this system specifies explicitly the powers of $t$ which depend on a choice of a sufficiently large $N^{*}\in\mathbb{N}$ and a sufficiently close to $v_{0}$ rational number $r^{*}<v_{0}$. More precisely, if $q:=p^{N^{*}}$ then $r^{*}(q-1):=b^{*}$ is a natural number prime to $p$, and there are no ramification breaks of the ramification filtration $\mathcal{G}^{(v)}_{<p,M}$ inside the interval $(r^{*},v_{0})$. (The ramification breaks of the filtration $\mathcal{G}^{(v)}_{<p,M}$ are discrete because they are discrete for abelian extensions and $\mathcal{G}_{<p,M}$ has finite nilpotence class.) Then the admissible exponents $t^{-\iota}$, $\iota\in\mathfrak{A}^{0}$, appear in the form $\iota=qp^{M-1}A-b^{*}U\in\mathfrak{A}^{0}$, where $A$ is a finite sum of various $p^{n}a$ with $n\geqslant 0$ and $a\in[0,(p-1)v_{0})$, and $U$ is a finite sum of $u_{j}p^{m_{j}}$ with $u_{j},m_{j}\geqslant 0$ and $1\leqslant\sum_{j}u_{j}<p$. One of properties of such exponents $\iota\in\mathfrak{A}^{0}$ is that $\mathrm{ch}(\iota):=\sum_{j}u_{j}$ is an invariant of $\iota$. In these terms, working with the Lie algebra $\bar{\mathcal{L}}$ inside $\bar{\mathcal{L}}^{{\dagger}}$ is related to $t$-exponents with $\mathrm{ch}(\iota)=0$, and working with $\mathrm{Ker}\bar{\mathcal{V}}\subset\bar{\mathcal{L}}^{{\dagger}}$ corresponds to t-exponents with $\mathrm{ch}(\iota)\geqslant 1$.
d) Let us come back to the identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$ from a) determined (up to conjugation) by a choice of a suitable $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$. Then the image $\bar{e}$ of $e$ under the projection $\mathcal{L}_{O_{M}(\mathcal{K})}\to\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ determines the epimorphic map $\bar{\eta}^{0}:{\mathcal{G}}_{<p,M}\to G(\bar{\mathcal{L}})$. On the other hand, choose an analog $\bar{e}^{{\dagger}}$ of $\bar{e}$ for $\bar{\mathcal{L}}^{{\dagger}}$ which is compatible with $\bar{e}$, i.e. such that $(\bar{\mathcal{V}}\otimes\mathrm{id})\bar{e}^{{\dagger}}=\bar{e}$. This will allow us to obtain a unique lift $\bar{\eta}^{{\dagger}}:\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}}^{{\dagger}})$ of $\bar{\eta}^{0}$ such that $\bar{\mathcal{V}}\bar{\eta}^{{\dagger}}=\bar{\eta}^{0}$.
e) Note that $\bar{\mathcal{L}}^{{\dagger}}$ also could be provided with a decreasing filtration of length $p$ such that the projection $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}$ is epimorphism of filtered Lie algebras. As a matter of fact, the filtration on $\bar{\mathcal{L}}^{{\dagger}}$ is defined in terms of invariant $\mathrm{ch}(\iota)\in[0,p-1]$. This allows us to define the action $\Omega$ of the elements $\gamma\in\mathbb{Z}/p^{M}$ on the system of coefficients $\mathfrak{A}^{0}$ by the correspondences $t^{-qp^{M-1}a}\mapsto t^{-qp^{M-1}a}E(t^{b^{*}})^{a\gamma p^{M-1}}$, where $E$ is the Artin-Hasse exponential.
f) We use the action $\Omega$ to obtain a deformation of the element $\bar{e}^{{\dagger}}$ to $\Omega({\bar{e}}^{{\dagger}})\in\bar{\mathcal{L}}^{{\dagger}}_{O_{M}(\mathcal{K})}$. This element determines the deformation $\Omega(\bar{\eta}^{{\dagger}})$ of $\bar{\eta}^{{\dagger}}$ which is given by a unipotent automorphism $A^{{\dagger}}$ of the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$. Equivalently, it appears from a differentiation $B^{{\dagger}}$ of $\bar{\mathcal{L}}^{{\dagger}}$ such that $A^{{\dagger}}=\widetilde{\exp}B^{{\dagger}}$ (this is the truncated exponential). With this notation, our main result states that the ramification ideal ¯$\bar{\mathcal{L}}^{(v_{0})}$ appears as $\bar{\mathcal{V}}B^{{\dagger}}(\mathrm{Ker}{\bar{\mathcal{V}}})$.
In the remaining part of the paper we are doing explicit calculations with specially chosen element $\bar{e}$ to find explicit generators of ¯$\bar{\mathcal{L}}^{(v_{0})}$ in the style of paper [3].
1. Lie algebra $\mathcal{L}$ and its filtration
As earlier, $\mathcal{K}$ is a complete discrete valuation field of characteristic $p$ with a finite residue field $k\simeq\mathbb{F}_{p^{N_{0}}}$ and a fixed uniformiser $t_{0}$, i.e. $\mathcal{K}=k((t_{0}))$. Denote by $\mathcal{K}_{sep}$ a separable closure of $\mathcal{K}$. Let $\mathcal{G}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K})$. Then $\mathcal{G}$ acts on $\mathcal{K}_{sep}$ and for any $a\in\mathcal{K}_{sep}$ and $g_{1},g_{2}\in\mathcal{G}$, $g_{1}(g_{2}a)=(g_{1}g_{2})a$. Denote by $\sigma$ the $p$-th power map on $\mathcal{K}$, $\mathcal{K}_{sep}$ and $k$.
1.1. Lie algebra $\mathcal{L}$
For any field extension $\mathcal{E}$ of $\mathcal{K}$ in $\mathcal{K}_{sep}$, we have a lift modulo $p^{M}$, $O_{M}(\mathcal{E}):=W_{M}(\sigma^{M-1}\mathcal{E})[t]\subset W_{M}(\mathcal{E})$, where $t=[t_{0}]\in W_{M}(\mathcal{K})$ is the Teichmuller representative of $t_{0}$, cf. e.g. [2]. The system of lifts $O_{M}(\mathcal{E})$ is functorial in $\mathcal{E}$, the actions of $\mathcal{G}$ and $\sigma$ can be naturally extended to $O_{M}(\mathcal{K}_{sep})$. Note that $O_{M}(\mathcal{K}_{sep})^{\mathcal{G}}=O_{M}(\mathcal{K})=W_{M}(k)((t))$ and $O_{M}(\mathcal{K}_{sep})|_{\sigma=\mathrm{id}}=W_{M}(\mathbb{F}_{p})$.
Let $\mathbb{Z}^{+}(p)=\{a\in\mathbb{N}\ |\ \mathrm{gcd}\,(a,p)=1\}$, $\mathbb{Z}^{0}(p)=\mathbb{Z}^{+}(p)\cup\{0\}$.
Let ^$\widehat{\mathcal{L}}_{W_{M}(k)}$ be a profinite free Lie $W_{M}(k)$-algebra with the (topological) module of (free) generators
(1.1) $\{D_{0}\}\cup\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p),n\in\mathbb{Z}/N_{0}\}\,.$
Define $\sigma$-action on ^$\widehat{\mathcal{L}}_{W_{M}(k)}$ by setting $\sigma:D_{an}\mapsto D_{a,n+1}$ and $\sigma:D_{0}\mapsto D_{0}$. Then ^$\widehat{\mathcal{L}}_{W_{M}(k)}|_{\sigma=\mathrm{id}}:=\widehat{\mathcal{L}}$ is a free profinite Lie $\mathbb{Z}_{p}/p^{M}$-algebra such that ^$\widehat{\mathcal{L}}\otimes W_{M}(k)=\widehat{\mathcal{L}}_{W_{M}(k)}$. We set :=^$\mathcal{L}:=\widehat{\mathcal{L}}/C_{p}(\widehat{\mathcal{L}})$. This is a maximal quotient of $\widehat{\mathcal{L}}$ of nilpotent class $<p$.
1.2. Ideals $\mathcal{L}(w)$, $w\geqslant 1$
Recall that we fixed $v_{0}\in\mathbb{R}_{>0}$.
Definition.
For $m\in\mathbb{Z}_{\geqslant 0}$ and $s\in\mathbb{N}$ , let $U(s,m)$ be the minimal value of $1/p^{i_{1}}+\ldots+1/p^{i_{s}}$, where all $i_{j}$ run over $\mathbb{Z}_{\geqslant 0}$ and satisfy the condition $i_{1}+\dots+i_{s}=m$. By definition, set for any $m\geqslant 0$, $U(0,m)=0$.
Definition.
For $w\geqslant 1$, let $\mathcal{L}(w)$ be the minimal central filtration of ideals in $\mathcal{L}$ such that if $a/v_{0}\geqslant U(w-1,m)$ then $p^{m}D_{an}\in\mathcal{L}(w)_{W_{M}(k)}$.
Note that $\mathcal{L}(1)=\mathcal{L}$ and $\forall w_{1},w_{2}\in\mathbb{N}$, $[\mathcal{L}(w_{1}),\mathcal{L}(w_{2})]\subset\mathcal{L}(w_{1}+w_{2})$.
Proposition 1.1.
Let $s,s^{\prime},m,m^{\prime}\in\mathbb{Z}_{\geqslant 0}$. Then:
a) $U(s,m)+U(s^{\prime},m^{\prime})\geqslant U(s+s^{\prime},m+m^{\prime})\,.$
b) For $s\geqslant 1$, let $m=su_{0}+s_{1}$, where $0\leqslant s_{1}<s$. Then
$U(s,m)=s_{0}/p^{u_{0}}+s_{1}/p^{u_{0}+1}\,\text{with }s_{0}=s-s_{1}\,.$
In particular, $u_{0},s_{0},s_{1}$ are uniquely determined by $m$ and $s$.
c) Suppose $s+s^{\prime}<p$. Then for any $M\geqslant 0$,
$\min\{U(s,m)+U(s^{\prime},m^{\prime})\ |\ m+m^{\prime}=M\}=U(s+s^{\prime},M)\,.$
Proof.
a) is obvious.
b) Suppose $m_{1},\dots,m_{s}\in\mathbb{Z}_{\geqslant 0}$, $m_{1}+\ldots+m_{s}=m$ and
$U(s,m)=1/p^{m_{1}}+\cdots+1/p^{m_{s}}\,.$
If for some indices $1\leqslant i,j\leqslant s$ it holds that $m_{i}\geqslant m_{j}+2$ then we obtain the contradiction by the use of the obvious inequality
$1/p^{m_{i}-1}+1/p^{m_{j}+1}<1/p^{m_{i}}+1/p^{m_{j}}\,.$
Therefore, the minimal value $U(s,m)$ can appear only if there is $u_{0}\in\mathbb{N}$ such that among all numbers $m_{i}$, $1\leqslant j\leqslant s$, there are $s_{0}>0$ of those which are equal to $u_{0}$ and all remaining $s_{1}=s-s_{0}$ numbers are equal to $u_{0}+1$. It remains to note that $m=m_{1}+\dots+m_{s}=u_{0}s_{0}+(u_{0}+1)s_{1}=su_{0}+s_{1}\,.$
c) Suppose $U(s,m)+U(s^{\prime},m^{\prime})$ is the minimal value of the LHS. We can assume that $s,s^{\prime}\geqslant 1$. Let $m=si+s_{1}$ and $m^{\prime}=s^{\prime}i^{\prime}+s_{1}^{\prime}$, where $0\leqslant s_{1}<s$ and $0\leqslant s_{1}^{\prime}<s^{\prime}$. Then for $s_{0}=s-s_{1}$ and $s_{0}^{\prime}=s^{\prime}-s_{1}^{\prime}$, $s_{0}+s_{1}+s_{0}^{\prime}+s_{1}^{\prime}<p$, $U(s,m)=s_{0}/p^{i}+s_{1}/p^{i+1}$ and $U(s^{\prime},m^{\prime})=s_{0}^{\prime}/p^{i^{\prime}}+s_{1}^{\prime}/p^{i^{\prime}+1}$.
$c_{1}$) Suppose $i=i^{\prime}$. Then $U(s,m)+U(s^{\prime},m^{\prime})=U(s+s^{\prime},m+m^{\prime})$ and our statement is proved.
$c_{2}$) Suppose $i>i^{\prime}$.
$c_{21}$) Let $s_{1}\neq 0$ and $s_{1}^{\prime}\neq s^{\prime}-1$. In this case $m-1=si+(s_{1}-1)$ and $m^{\prime}=s^{\prime}i^{\prime}+(s_{1}^{\prime}+1)$ imply that
(1.2) $U(s,m-1)+U(s^{\prime},m^{\prime}+1)<U(s,m)+U(s^{\prime},m^{\prime})$
because $U(s,m-1)=U(s,m)+(1/p^{i}-1/p^{i+1})$ and $U(s^{\prime},m^{\prime}+1)=U(s^{\prime},m^{\prime})-(1/p^{i^{\prime}}-1/p^{i^{\prime}+1})$.
$c_{22}$) Let $s_{1}=0$ but $s_{1}^{\prime}\neq s^{\prime}-1$. In this case $m-1=s(i-1)+(s-1)$ and $m+1=s^{\prime}i^{\prime}+(s^{\prime}_{1}+1)$ imply that either we have again the contradiction (1.2) or $i=i^{\prime}+1$. But in the later case, $m+m^{\prime}=(s+s^{\prime})i^{\prime}+s+s_{1}^{\prime}$ and
$U(s+s^{\prime},m+m^{\prime})=\frac{s^{\prime}}{p^{i^{\prime}}}+\frac{s+s_{1}^{\prime}}{p^{i^{\prime}+1}}=U(s,m)+U(s^{\prime},m^{\prime})\,.$
$c_{23}$) Let $s_{1}\neq 0$ but $s_{1}^{\prime}-1=s^{\prime}$. Here $m-1=si+s_{1}-1$ and $m+1=(s^{\prime}+1)i^{\prime}$. Then $U(s,m-1)=U(s,m)+(1/p^{i}-1/p^{i+1})$ and $U(s^{\prime},m^{\prime}+1)=U(s^{\prime},m^{\prime})-(1/p^{i^{\prime}}-1/p^{i^{\prime}+1})$. Then again we obtain contradiction (1.2) because $i>i^{\prime}$.
$c_{24}$) Let $s_{1}=0$ and $s_{1}^{\prime}=s^{\prime}-1$. Here $m-1=s(i-1)+s-1$ and $m^{\prime}+1=s^{\prime}(i^{\prime}+1)$. We obtain either contradiction (1.2) or $i-1=i^{\prime}$. In the later case $m+m^{\prime}=s^{\prime}i^{\prime}+s+s^{\prime}-1$ and $U(s+s^{\prime},m+m^{\prime})=U(s,m)+U(s^{\prime},m^{\prime})$. ∎
Example. Suppose $m=0$ and $s\geqslant 1$. Then $u_{0}=s_{1}=0$ and for all $w\geqslant 1$, $U(w-1,0)=w-1$. If $(w-1)v_{0}\leqslant a<wv_{0}$ then $D_{an}\in\mathcal{L}(w)_{W_{M}(k)}$. In particular, if $a\geqslant(p-1)v_{0}$ then $D_{an}\in\mathcal{L}(p)_{W_{M}(k)}$. Notice that in the modulo $p$ situation (i.e. when $M=1$) from [12] the ideals $\mathcal{L}(w)$ coincide with the ideals introduced in that paper (and denoted by $\mathcal{L}(s)$).
Proposition 1.1b) means:
Proposition 1.2.
The filtration $\{\mathcal{L}(w)\ |\ w\geqslant 1\}$ is the minimal central filtration of $\mathcal{L}$ such that for all $a\in\mathbb{Z}^{0}(p)$ and non-negative integers $\alpha_{0}\neq 0,\alpha_{1},j$, hold the conditions
$\bullet$ $R_{a}(\alpha_{0},\alpha_{1},j)$: if $a/v_{0}\geqslant\alpha_{0}/p^{j}+\alpha_{1}/p^{j+1}$ then
$p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(1+\alpha_{0}+\alpha_{1})_{W_{M}(k)}\,.$
1.3. Independent relations $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$, $s<p$
Definition.
Given $a\in\mathbb{Z}^{0}(p)$ and $0\leqslant s<p$, let $M(a,s)$ be the minimal non-negative integer such that $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$.
Remark.
The restriction $s<p$ is introduced because we are going to work below with the Lie algebra ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$ provided with the filtration by the ideals ¯$\bar{\mathcal{L}}(w)=\mathcal{L}(w)/\mathcal{L}(p)$.
Clearly, $M(a,0)=0$ and for all $n$, $p^{m}D_{0}\in\mathcal{L}(1)_{W_{M}(k)}\setminus\mathcal{L}(2)_{W_{M}(k)}$. In addition, if $a/v_{0}\geqslant U(p-1,0)=p-1$ then $D_{an}\in\mathcal{L}(p)_{W_{M}(k)}$. Therefore, we can assume below that $s\geqslant 1$ and $a<pv_{0}$.
Suppose $a/v_{0}=\sum_{i\geqslant 0}\theta_{i}(a)/p^{i}$ is the standard decomposition with the $p$-digits $0\leqslant\theta_{i}(a)<p$. Let $u(a)$ be the minimal index such that $\theta_{u(a)}(a)\neq 0$.
Set $u=u(a)$, $\kappa_{0}=\theta_{u}(a)$ and $\kappa_{1}=\theta_{u+1}(a)$. Note that
$\kappa_{0}/p^{u}+\kappa_{1}/p^{u+1}\leqslant a/v_{0}<\kappa_{0}/p^{u}+(\kappa_{1}+1)/p^{u+1}\,.$
Consider the relation $R_{a}(\alpha_{0},\alpha_{1},j)$ from Prop.1.2. Then
(1.3) $\alpha_{0}/p^{j}+\alpha_{1}/p^{j+1}\leqslant\kappa_{0}/p^{u}+\kappa_{1}/p^{u+1}\,,$
and $p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(1+\alpha_{0}+\alpha_{1})_{W_{M}(k)}\,.$
Note that $\alpha_{0}+\alpha_{1}\geqslant p-1$ implies $p^{\alpha_{0}j+\alpha_{1}(j+1)}D_{an}\in\mathcal{L}(p)_{W_{M}(k)}$. If we want to obtain the relations for the ideals $\mathcal{L}(s)$ modulo $\mathcal{L}(p)$ then it makes sense to restrict ourselves with the cases where $\alpha_{0}+\alpha_{1}<p$. In particular, $\alpha_{0}/p^{j}+\alpha_{1}/p^{j+1}$ will appear as a standard decomposition in base $p$ with $p$-digits $\alpha_{0},\alpha_{1}$. Now inequality (1.3) implies one of the following options:
$j\geqslant u+1$;
$j=u$ and $\alpha_{0}<\kappa_{0}$;
$j=u$, $\alpha_{0}=\kappa_{0}$ and $\alpha_{1}\leqslant\kappa_{1}$.
Proceeding along these lines we obtain the following list of relations of the form $p^{M(a,s)}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$:
Proposition 1.3.
For $(\alpha_{0},\alpha_{1},u)$ such that:
a) $1\leqslant\alpha_{0}\leqslant\kappa_{0}$ and $\alpha_{1}=0$, we have $s=\alpha_{0}$ and
$M(a,s)=\alpha_{0}u\,;$
b) $\alpha_{0}=\kappa_{0}$ and $1\leqslant\alpha_{1}\leqslant\kappa_{1}$, we have $s=\kappa_{0}+\alpha_{1}$ and
$M(a,s)=M(a,\kappa_{0})+\alpha_{1}(u+1)\,;$
c) $\alpha_{0}=\kappa_{0}-1$, $\alpha_{1}=\kappa_{1}+2$, we have $s=\kappa_{0}+\kappa_{1}+1$ and
$M(a,s)=M(a,\kappa_{0}+\kappa_{1})+u+2\,;$
d) $\alpha_{0}=\kappa_{0}-1$ and $\alpha_{1}=\kappa_{1}+2+t<p$ with $t\geqslant 1$, we have
$s=\kappa_{0}+\alpha_{1}$ and
$M(a,s)=M(a,\kappa_{0}+\kappa_{1}+1)+t(u+1)\,.$
Remark.
Taking into account that we need relations modulo $\mathcal{L}(p)_{k}$:
– for $s\in[0,\kappa_{0}]$, $M(a,s)$ is increasing in $s$ with step $u$;
– for $s\in[\kappa_{0},\min\{\kappa_{0}+\kappa_{1},p-1\}]$ the step is $u+1$;
– for $s\in[\kappa_{0}+\kappa_{1},\min\{\kappa_{0}+\kappa_{1}+1,p-1\}]$ the step is $u+2$;
– if $s\in[\kappa_{0}+\kappa_{1}+1,p-1]$ the step is $u+1$.
Proposition 1.4.
Relations from Proposition 1.3 with $s<p$ are all independent relations of the form $p^{A}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$, where $A\geqslant 0$ and $10\leqslant s\leqslant p-$ 1.
Proof.
The independence is obvious because for each $a$, the numbers $M(a,s)$ are strictly increasing in $s$.
Let $a/v_{0}\geqslant s_{0}/p^{j}+s_{1}/p^{j+1}$ with integers $j\geqslant 0$, $s_{0}\geqslant 1$ and $s_{1}\geqslant 0$. Note that $s_{0}$ and $s_{1}$ are not necessarily $p$-digits.
Prove that the corresponding condition $p^{m}D_{an}\in\mathcal{L}(1+s)_{W_{M}(k)}$, where $m=s_{0}j+s_{1}(j+1)$ and $s=\min\{s_{0}+s_{1},p-1\}$, is implied by one of relations listed in Proposition 1.3.
Otherwise, there is at least one triple $(\alpha_{0},\alpha_{1},u)$ listed in Prop. 1.3 such that the following two conditions hold
(1.4) $\alpha_{0}+\alpha_{1}\leqslant s_{0}+s_{1}$
(1.5) $\alpha_{0}u+\alpha_{1}(u+1)>s_{0}j+s_{1}(j+1)\,.$
Then $s_{0}\geqslant s_{0}-\alpha_{0}\geqslant(\alpha_{0}+\alpha_{1}-s_{0}-s_{1})(u+1)+(s_{0}-\alpha_{0})=\alpha_{0}u+\alpha_{1}(u+1)-(s_{0}u+s_{1}(u+1))>(s_{0}+s_{1})(j-u)$ implies that $j\leqslant u$.
If $j<u$ then $p/p^{u}>a/v_{0}\geqslant s_{0}/p^{j}$ implies $p>s_{0}p^{u-j}\geqslant p$. The contradiction.
Therefore, $j=u$. Then (1.4) implies $\alpha_{1}>s_{1}$ because
$0\geqslant(\alpha_{0}+\alpha_{1}-s_{0}-s_{1})u>s_{1}-\alpha_{1}\,.$
In particular, $\alpha_{1}\neq 0$ and the couple $(\alpha_{0},\alpha_{1})$ does not come from condition a).
The condition $\alpha_{0}+\alpha_{1}\leqslant s_{0}+s_{1}$ also implies that $\alpha_{0}<s_{0}$. This excludes the case b) because $s_{0}\leqslant\kappa_{0}$ (use that $j=u$).
It remains to consider the cases c) and d), where $\alpha_{0}=\kappa_{0}-1$ and $s_{0}=\kappa_{0}$. Here $\alpha_{0}=\kappa_{0}-1$, $\alpha_{1}\geqslant\kappa_{1}+2$, $s_{0}=\alpha_{0}+1$, and $s_{1}\leqslant\kappa_{1}$. Then we obtain the contradiction $\alpha_{0}+\alpha_{1}\geqslant\kappa_{0}+\kappa_{1}+1>s_{0}+s_{1}$. ∎
1.4. Explicit description of ideals $\mathcal{L}(w)_{W_{M}(k)}$
Recall that the ideal $\mathcal{L}(w)_{W_{M}(k)}$ is generated as $W_{M}(k)$-module by all commutators
$p^{m}[\dots[D_{a_{1}n_{1}},D_{a_{2}n_{2}}],\dots,D_{a_{r}n_{r}}]\,:=p^{m}[D_{\bar{a}\bar{n}}]$
where $\bar{a}=(a_{1},\ldots,a_{r})$, for $1\leqslant i\leqslant r:=r(\bar{a})$, there are $m_{i},s_{i}\geqslant 0$ such that $m=\sum_{i}m_{i}$,  $r+\sum_{i}s_{i}\geqslant w$ and all $a_{i}\geqslant U(s_{i},m_{i})v_{0}$.
Definition.
Set $M(\bar{a},s)=\min\{m\in\mathbb{Z}_{\geqslant 0}\ |\ p^{m}[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r(\bar{a})+s)_{W_{M}(k)}\}$.
We always assume that $r=r(\bar{a})<p$ because $D_{\bar{a}\bar{n}}\in\mathcal{L}(r)_{W_{M}(k)}$.
For $1\leqslant i\leqslant r$, consider the following decompositions in the “base $p$” with $p$-digits from $[0,p)$:
[…]
Set $\bar{u}=u(\bar{a})=\min\{\ j\ |\ (\theta_{j}(a_{1}),\ldots,\theta_{j}(a_{r}))\neq\bar{0}\}$.
Set $\bar{\kappa}_{0}=\kappa_{0}(\bar{a})=\sum_{1\leqslant i\leqslant r}\theta_{\bar{u}}(a_{i})$ and $\bar{\kappa}_{1}=\kappa_{1}(\bar{a})=\sum_{1\leqslant i\leqslant r}\theta_{\bar{u}+1}(a_{i})$.
Proposition 1.5.
a) $M(\bar{a},j)=j\bar{u}$ for $0\leqslant j\leqslant\min\{p-r,\bar{\kappa}_{0}\}$;
b) $M(\bar{a},\bar{\kappa}_{0}+j)=\bar{u}\bar{\kappa}_{0}+(\bar{u}+1)j$, for $1\leqslant j\leqslant\min\{p-(r+\bar{\kappa}_{0}),\bar{\kappa}_{1}\}$;
c) $M(\bar{a},\bar{\kappa}_{0}+\bar{\kappa}_{1}+j)=\bar{\kappa}_{0}\bar{u}+\bar{\kappa}_{1}(\bar{u}+1)+1+j(\bar{u}+1)$, where $j$ runs over $1\leqslant j\leqslant p-(r+\bar{\kappa}_{0}+\bar{\kappa}_{1})$.
Remark.
If $\bar{u}=0$ then $[D_{\bar{a}\bar{n}}]\in\mathcal{L}(r+\bar{\kappa}_{0})_{W_{M}(k)}$.
1.5. $\mathbb{Z}_{p}$-submodules $\mathcal{L}_{\mathcal{N}}(w)\subset\mathcal{L}(w)$
Definition.
For $w\geqslant 1$, let $\mathcal{L}_{\mathcal{N}}(w)$ be the closed submodule in $\mathcal{L}$ such that $\mathcal{L}_{\mathcal{N}}(w)_{W_{M}(k)}$ is generated by all $p^{m}[D_{\bar{a}\bar{n}}]$ which satisfy the inequality $|\bar{a}|\geqslant v_{0}U(w-1,m)$.
Proposition 1.6.
We have the following decomposition of $\mathbb{Z}_{p}$-modules
$\mathcal{L}_{\mathcal{N}}(w)\cap C_{2}(\mathcal{L})\subset\sum_{w_{1}+w_{2}\geqslant w}[\mathcal{L}_{\mathcal{N}}(w_{1}),\mathcal{L}_{\mathcal{N}}(w_{2})]$
It will be sufficient to prove the following lemma.
Lemma 1.7.
If $p^{m}[D_{\bar{a}_{1}\bar{n}_{1}},D_{\bar{a}_{2}\bar{n}_{2}}]\in\mathcal{L}_{\mathcal{N}}(w)_{W_{M}(k)}$ then for $i=1,2$, there are $m_{i},w_{i}$ such that:
1)  $m_{1}+m_{2}=m$, $w_{1}+w_{2}=w$;
2)  $p^{m_{i}}[D_{\bar{a}_{i}\bar{n}_{i}}]\in\mathcal{L}_{\mathcal{N}}(w_{i})_{W_{M}(k)}$.
Proof of Lemma.
$\bullet$ Choose $w_{1},w_{2}\geqslant 1$ with maximal $w_{1}+w_{2}$ s.t. for some $m_{1},m_{2}$, $m=m_{1}+m_{2}$ and for $i=1,2$, $|\bar{a}_{i}|\geqslant v_{0}U(w_{i}-1,m_{i})\,.$
We must prove that $w_{1}+w_{2}\geqslant w$.
Suppose $w_{1}+w_{2}<w$.
We can assume that (otherwise, $w_{1}+w_{2}$ is not maximal) for $i=1,2$,
$v_{0}U(w_{i},m_{i})>|\bar{a}_{i}|\geqslant v_{0}U(w_{i}-1,m_{i})\,.$
Let $v_{0}U(w_{1}+w_{2},m)>|\bar{a}_{1}|+|\bar{a}_{2}|$. Then $U(w_{1}+w_{2},m)>U(w-1,m)$ implies $w_{1}+w_{2}\geqslant w$, q.e.d.
Otherwise, $|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant v_{0}U(w_{1}+w_{2},m)$. By Prop. 1.1c), $\exists m_{1}^{\prime},m_{2}^{\prime}$ such that  $m_{1}^{\prime}+m_{2}^{\prime}=m$\&& $U(w_{1},m_{1}^{\prime})+U(w_{2},m_{2}^{\prime})=U(w_{1}+w_{2},m)$.
Let $m_{1}^{\prime}\geqslant m_{1}$ and $m_{2}^{\prime}\leqslant m_{2}$. Then $v_{0}U(w_{1},m_{1}^{\prime})\geqslant v_{0}U(w_{1},m_{1})>|\bar{a}_{1}|$ and $v_{0}U(w_{2},m_{2}^{\prime})\leqslant v_{0}U(w_{2},m_{2})$.
If $|\bar{a}_{2}|<v_{0}U(w_{2},m^{\prime}_{2})$ then $v_{0}U(w_{1}+w_{2},m)>|\bar{a}_{1}|+|\bar{a}_{2}|$. The contradiction.
If $|\bar{a}_{2}|\geqslant v_{0}U(w_{2},m_{2}^{\prime})$ then $w_{2}$ can be replaced with a bigger one and $w_{1}+w_{2}$ is not maximal. The lemma is proved. ∎
Corollary 1.8.
$\mathcal{L}_{\mathcal{N}}(w)\subset\mathcal{L}(w)$.
Proof.
By induction on $w$ it will be enough to verify the statement for $p^{m}[D_{\bar{a}\bar{n}}]$ with $r(\bar{a})=1$. But this follows from the definition of $\mathcal{L}(w)$. ∎
Definition.
Let ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$ with the induced filtration by the ideals ¯$\bar{\mathcal{L}}(w)=\mathcal{L}(w)/\mathcal{L}(p)$, $w\geqslant 1$.
Clearly, ¯$\bar{\mathcal{L}}(p)=0$ and ¯$\bar{\mathcal{L}}_{W_{M}(k)}$ is generated by $D_{0}$ and all $D_{an}$ with $a<(p-1)v_{0}$.
2. Lie algebra $\bar{\mathcal{L}}^{\,{\dagger}}$ and ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}\subset\bar{\mathcal{L}}$
2.1. Parameters $r^{*}$ and $N^{*}$
Let $u^{*}=(p-1)(p-2)+1$ and $w^{*}=(p-1)v_{0}$.
For $1\leqslant u\leqslant u^{*}$, introduce the following subset in $\mathbb{Q}$ :
$A[u]$ is the set of all
$a_{1}p^{n_{1}}+a_{2}p^{n_{2}}+\dots+a_{u}p^{n_{u}}\,,$
where $n_{1},\dots,n_{u}\leqslant 0$, $n_{1}=0$ and all $a_{i}\in[0,w^{*})\cap\mathbb{Z}$. For $m\in\mathbb{Z}_{\geqslant 0}$, denote by $A[u,m]$ the subset of $A[u]$ consisted of the elements such that all $n_{u}\geqslant-m$.
For $1\leqslant s<p$, $0\leqslant m<M$, let $\delta_{0}(s,m)$ be the minimum of positive values of $v_{0}-\alpha/U(s,m)$ where $\alpha\in A[u,m]$ and $u\leqslant u^{*}$. The existence of such $\delta_{0}(s,m)$ can be proved easily by induction on $u$ for any fixed $s$ and $m$.
[…]
Let $r^{*}\in\mathbb{Q}$ be such that $r^{*}=b_{0}^{*}/(q^{*}_{0}-1)$, where $q^{*}_{0}=p^{N^{*}_{0}}$ with $N^{*}_{0}\geqslant 2$, $b^{*}_{0}\in\mathbb{N}$ and $\mathrm{gcd}(b_{0}^{*},\,p(q^{*}_{0}-1))=1$. The set of such $r^{*}$ is dense in $\mathbb{R}_{>0}$ and we can assume that $r^{*}\in(v_{0}-\delta_{0},v_{0})$.
For $1\leqslant u\leqslant u^{*}$, introduce the following subset in $\mathbb{Q}$ :
$B[u]$ is the set of all numbers
$r^{*}(b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}})\,,$
where all $m_{1},\ldots,m_{u}\leqslant M-1$, $m_{1}\geqslant 0$, $b_{i}\geqslant 0$, $b_{1}\neq 0$, $b_{1}+\dots+b_{u}<p$. (In particular, $0\notin B[u]$.) For $m\in\mathbb{Z}_{\geqslant 0}$, $B[u,m]$ is the subset of $B[u]$ consisted of the elements such that all $m_{u}\geqslant-m$.
The above sets $A[u]$ and an analog of the set $B_{1}[u]:=p^{-(M-1)}B[u]$ (where we required $m_{1}=0$ instead of $-M<m_{1}\leqslant 0$) were introduced in [12]. Proceeding similarly to [12] we obtain that for any $u\leqslant u^{*}$, $A[u]\cap B_{1}[u]=\emptyset$ and there is $\rho(u^{*})>0$ such that
$\rho(A[u],B_{1}[u]):=\inf\{|\alpha-\beta|\ |\ \alpha\in A[u],\beta\in B_{1}[u]\}\geqslant\rho(u^{*})\,.$
Choose $N^{*}\in\mathbb{N}$ satisfying the following conditions:
C1)  $N^{*}\equiv 0\,\mathrm{mod}\,N^{*}_{0}$;
C2)  $p^{N^{*}}\rho(A[u^{*}],B_{1}[u^{*}])\geqslant 2r^{*}(p-1)$;
C3)  $r^{*}(1-p^{-N^{*}})\in(v_{0}-\delta_{0},v_{0})$.
Introduce $q=p^{N^{*}}$ and $b^{*}=b^{*}_{0}(q-1)/(q_{0}-1)\in\mathbb{N}$.
Note that $r^{*}=b^{*}/(q-1)$ and $b^{*}\in\mathbb{Z}^{+}(p)$.
Proposition 2.1.
a) If $\alpha\in A[u^{*}]$ and $\beta\in B[u^{*}]$ then
$q\,|qp^{M-1}\alpha-(q-1)\beta|>p^{M-1}b^{*}(p-1)\,;$
b) If $\alpha\in A[u^{*}]$ and $\alpha/U(s,m)<v_{0}$ then $q\alpha-b^{*}U(s,m)<0$.
Proof.
a) Use condition C2). Let $\beta_{1}=\beta/p^{M-1}\in B_{1}[u^{*}]$. Then
$q\,|q\alpha-(q-1)\beta_{1}|=q^{2}|\alpha-\beta_{1}+\beta_{1}/q|\geqslant q^{2}|\alpha-\beta_{1}|-\beta q\geqslant q^{2}\rho(A[u^{*}],\,B_{1}[u^{*}])$
$-r^{*}(p-1)q\geqslant 2r^{*}(p-1)q-r^{*}(p-1)q=r^{*}(p-1)q>b^{*}(p-1)\,.$
It remains to multiply this inequality by $p^{M-1}$;
b) Use condition C3): $\alpha/U(s,m)<v_{0}$ implies that $\alpha/U(s,m)<v_{0}-\delta_{0}$ and $q\alpha/U(s,m)<q(v_{0}-\delta_{0})<qr^{*}(1-q^{-1})=b^{*}$.
2.2. The set $\mathfrak{A}^{0}$
Use the above parameters $r^{*}$, $N^{*}$, $q=p^{N^{*}}$.
Definition.
${\mathfrak{A}}^{0}$ is the set of all $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)$, where $m\geqslant 0$, $\alpha\in A[u^{*},m]$, $\beta\in B[u^{*},m]\cup\{0\}$ and $|\iota|\leqslant p^{M-1}b^{*}(p-1)$. (Note that $p^{m}\alpha\in\mathbb{Z}_{\geqslant 0}$ and $p^{m}\beta/r^{*}\in\mathbb{N}$.)
Remark.
a) Prop.2.1 implies that for all $\iota\in\mathfrak{A}^{0}$, we have $m<N^{*}$.
b) Any $p^{m}\alpha$ appears in the form $a_{1}p^{n_{1}}+\ldots+a_{u^{*}}p^{n_{u^{*}}}$, where all $a_{i}\in[0,(p-1)v_{0})$ and $0\leqslant n_{i}<N^{*}$.
c) All $p^{m}\beta$ appear in the form $b^{*}(b_{1}p^{m_{1}}+\ldots+b_{u^{*}}p^{m_{u^{*}}})$, where all $m_{i}\geqslant 0$ and $11\leqslant b_{1}+\ldots+b_{u^{*}}\leqslant p-$ 1.
Let $\mathfrak{A}^{0}_{0}:=\{\iota\in\mathfrak{A}^{0}\ |\ \beta=0\}$.
Lemma 2.2.
Suppose $\iota=p^{m}(q\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$. Then:
a) $\mathfrak{A}^{0}_{0}=\{qp^{M-1}a\ |\ a\in[0,(p-1)v_{0})\cap\mathbb{Z}\,\}$;
b) if $\beta\neq 0$ then $m<N^{*}$ (in particular, $\mathfrak{A}^{0}$ is finite);
c) the integers $p^{m}\alpha$ and $p^{m}\beta/r^{*}$ do not depend on the presentation of $\iota$ in the form $p^{m}(qp^{M-1}\alpha-(q-1)\beta)$ from the definition of $\mathfrak{A}^{0}$.
Proof.
a) If $\iota\in\mathfrak{A}^{0}_{0}$ then $\iota=qp^{M-1}(p^{m}\alpha)\in\mathfrak{A}^{0}_{0}\subset\mathfrak{A}^{0}$ means that
$p^{m}\alpha/(p-1)\leqslant b^{*}/q=r^{*}(1-q^{-1})\in(v_{0}-\delta_{0},v_{0})\,.$
Note that $U(p-1,0)=p-1$ and $\delta_{0}\leqslant\delta(p-1,0)$, cf. Sect.2.1. Threfore, the inequalities $p^{m}\alpha/(p-1)<v_{0}$ and $p^{m}\alpha/(p-1)\leqslant v_{0}-\delta_{0}$ are equivalent. As a result,
$\mathfrak{A}^{0}_{0}\subset\{qp^{M-1}a\ |\ a\in[0,(p-1)v_{0})\cap\mathbb{Z}\}\,.$
The opposite embedding is obvious.
b) If $m\geqslant N^{*}$ then by Prop. 2.1, $|\iota|>p^{M-1}b^{*}(p-1)$ i.e. $\iota\notin\mathfrak{A}^{0}$.
c) If $\iota=p^{m^{\prime}}(qp^{M-1}\alpha^{\prime}-(q-1)\beta^{\prime})$ is another presentation of $\iota$ then $p^{m}\beta/r^{*}$ and $p^{m^{\prime}}\beta^{\prime}/r^{*}$ are non-negative congruent modulo $qp^{M-1}$ integers and the both are smaller than $qp^{M-1}$. Indeed, we have
$\beta/r^{*}=b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}}\,,$
where all $m_{i}\leqslant M-1$ and $u\leqslant u^{*}$, then
$p^{m}\beta/r^{*}\leqslant p^{m+M-1}(b_{1}+\dots+b_{u})\leqslant p^{m+M-1}(p-1)<p^{m+M}\leqslant qp^{M-1}$
because $m<N^{*}$. Similarly, $p^{m^{\prime}}\beta^{\prime}/r^{*}<qp^{M-1}$. Therefore, they coincide and this implies also that $p^{m}\alpha=p^{m^{\prime}}\alpha^{\prime}$. ∎
Corollary 2.3.
Suppose that $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$. Then the sum of the “ $p$-digits” $b_{1}+\dots+b_{u}$ of the appropriate $\beta/r^{*}=b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}}$ depends only on $\iota$.
Definition.
$\mathrm{ch}\,\iota:=b_{1}+\dots+b_{u}$
Remark.
If $\mathrm{ch}\,\iota\geqslant 1$ then $\iota\neq 0$. (Use that $A[u]\cap B_{1}[u]=\emptyset$.)
In the notation from Sect. 2.2 let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in{\mathfrak{A}}^{0}$. By Lemma 2.2, $p^{m}\alpha$ depends only on $\iota$ and can be presented (non-uniquely) in the form $a_{1}p^{n_{1}}+a_{2}p^{n_{2}}+\dots+a_{u}p^{n_{u}}$ where all coefficients $a_{i}\in[0,(p-1)v_{0}))\cap\mathbb{Z}$, $0\leqslant n_{i}\leqslant m$, $n_{1}=m$ and $u\leqslant u^{*}$.
Definition.
$\kappa(\iota)$ is the maximal natural number such that for any above presentation of $p^{m}\alpha$, $\kappa(\iota)\leqslant u$.
Remark.
a) If $\iota\in\mathfrak{A}^{0}$ then $\kappa(\iota)\leqslant u^{*}$ and $\mathrm{ch}(\iota)\leqslant p-1$;
b) if $\iota\in\mathfrak{A}^{0}_{0}$ then $\mathrm{ch}(\iota)=0$;
c) if $\iota\in\mathfrak{A}^{0}_{0}$ and $\iota\neq 0$ then $\kappa(\iota)=1$.
2.3. Lie algebras ${\mathcal{L}}^{\,{\dagger}}$ and $\bar{\mathcal{L}}^{\,{\dagger}}$
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ be given in the standard notation from Sect. 2.2.
Definition.
The subset $\mathfrak{A}^{+}(p)$ consists of $\iota\in\mathfrak{A}^{0}$ such that
$\iota>0$;
$\mathrm{gcd}(p^{m}\alpha\,,p^{m}\beta/r^{*},\,p)=1$;
$\kappa(\iota)\leqslant(p-2)\mathrm{ch}(\iota)+1$.
Remark.
For $\iota\in\mathfrak{A}^{+}(p)$, $(p-2)\mathrm{ch}(\iota)+1\leqslant(p-2)(p-1)+1=u^{*}\,.$
The elements of $\{t^{-\iota}\ |\ \iota\in\mathfrak{A}^{+}(p)\}$ behave “well” modulo $(\sigma-\mathrm{id})\mathcal{K}$, i.e. the natural map $\sum\limits_{\iota\in\mathfrak{A}^{+}(p)}kt^{-\iota}\longrightarrow\mathcal{K}/(\sigma-\mathrm{id})\mathcal{K}$ is injective. This is implied by the following proposition.
Proposition 2.4.
Let $v_{p}$ be the $p$-adic valuation such that $v_{p}(p)=1$.
a) Then all $\iota p^{-v_{p}(\iota)}$, where $\iota\in\mathfrak{A}^{+}(p)$, are pairwise different.
b) If $\iota\in\mathfrak{A}^{+}(p)$ and $\mathrm{ch}(\iota)=1$ then $\iota p^{-v_{p}(\iota)}\geqslant qv_{0}-b^{*}$.
Proof.
a) Suppose $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{+}(p)$.
If $\mathrm{ch}(\iota)=0$ then $\iota\mapsto\iota p^{-v_{p}(\iota)}$ identifies $\{\iota\in\mathfrak{A}^{+}(p)\ |\ \mathrm{ch}(\iota)=0\}$ with $\mathbb{Z}^{+}(p)\cap[0,(p-1)v_{0})$, cf. Lemma 2.2a).
If $\mathrm{ch}(\iota)\geqslant 1$ then $\iota p^{-m}\notin p^{M}\mathbb{N}$, i.e. $m+M>v_{p}(\iota)$.
Indeed, $\iota p^{-m}=qp^{M-1}\alpha-(q-1)\beta\in p^{M}\mathbb{N}$ implies (use that $qp^{M-1}\alpha\in p^{M}\mathbb{N}$ because $m<N^{*}$) that $\beta=p^{-m}(b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}})\in p^{M}\mathbb{N}$ with all $m_{i}\leqslant m+M-1$. But this number is $\leqslant p^{M-1}\mathrm{ch}\,\iota<p^{M}$. The contradiction. Then Prop. 2.1a) implies that
$\iota p^{-v_{p}(\iota)}\geqslant\iota p^{-m-M+1}=p^{-M+1}|qp^{M-1}\alpha-(q-1)\beta|>$
$(b^{*}/q)(p-1)=r^{*}(1-q^{-1})(p-1)>(v_{0}-\delta_{0})(p-1)$
(use property C3 from Sect. 2.1).
Finally, if $\iota\in\mathfrak{A}^{0}_{0}$ then $\iota p^{-v_{p}(\iota)}=a<(p-1)v_{0}$ implies that $a<(v_{0}-\delta_{0})(p-1)$ by the choice of $\delta_{0}$, cf. Sect. 2.1. On the other hand, for all $\iota\in\mathfrak{A}^{+}(p)$ with $\mathrm{ch}(\iota)\geqslant 1$, the values $\iota p^{-v_{p}(\iota)}$ are different (use that $\mathrm{gcd}(p^{m}\alpha,p^{m}\beta/r^{*})\not\equiv 0\,\mathrm{mod}\,p$) and are bigger than $(v_{0}-\delta_{0})(p-1)$.
b) There is $10\leqslant A\leqslant M-$ 1 such that $\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})$, i.e. $\iota p^{-v_{p}(\iota)}=q(p^{A}\alpha)-b^{*}$ and b) holds when $p^{A}\alpha\geqslant v_{0}$. But if $p^{A}\alpha=\alpha/U(1,A)<v_{0}$ then Prop.2.1b) implies that $q\alpha-b^{*}U(1,A)=p^{-A}(q(p^{A}\alpha)-b^{*})<0$. The contradiction.
The proposition is completely proved. ∎
Definition.
$\mathfrak{A}^{0}(p)=\mathfrak{A}^{+}(p)\cup\{0\}$.
Let ^$\widehat{\mathcal{L}}_{W_{M}(k)}^{\,{\dagger}}$ be the Lie algebra over $W_{M}(k)$ with the set of free generators
$\{\ D^{\,{\dagger}}_{\iota n}\ |\ \iota\in\mathfrak{A}^{+}(p),n\in\mathbb{Z}/N_{0}\}\cup\{D^{\,{\dagger}}_{0}\}\,.$
Use notation $\sigma$ for the $\sigma$-linear automorphism of ^$\widehat{\mathcal{L}}^{\,{\dagger}}_{W_{M}(k)}$ such that $\sigma:D^{\,{\dagger}}_{\iota n}\mapsto D^{\,{\dagger}}_{\iota,n+1}$ and $\sigma:D^{{\dagger}}_{0}\mapsto D^{{\dagger}}_{0}$. Introduce the Lie $\mathbb{Z}/p^{M}$-algebras ^:=^$\widehat{\mathcal{L}}^{\,{\dagger}}:=\widehat{\mathcal{L}}^{\,{\dagger}}_{W_{M}(k)}|_{\sigma=\mathrm{id}}$ and =^$\mathcal{L}^{\,{\dagger}}=\widehat{\mathcal{L}}^{\,{\dagger}}/C_{p}(\widehat{\mathcal{L}}^{\,{\dagger}})$. Note that ^$\widehat{\mathcal{L}}^{\,{\dagger}}\otimes W_{M}(k)=\widehat{\mathcal{L}}^{\,{\dagger}}_{W_{M}(k)}$, and this matches the agreement about extensions of scalars from the Introduction.
Denote by $\{\mathcal{L}^{\,{\dagger}}(w)\}_{w\geqslant 1}$ the minimal central filtration of $\mathcal{L}^{\,{\dagger}}$ such that $p^{m}D^{\,{\dagger}}_{\iota n}\in\mathcal{L}^{\,{\dagger}}(w)_{W_{M}(k)}$ if $\iota\geqslant p^{M-1}b^{*}U(w-1,m)$.
This means that $\mathcal{L}^{\,{\dagger}}(w)_{W_{M}(k)}$ is an ideal in $\mathcal{L}^{\,{\dagger}}_{W_{M}(k)}$ generated as $W_{M}(k)$-module by all $p^{m}[\dots[D^{\,{\dagger}}_{\iota_{1}n_{1}},D^{\,{\dagger}}_{\iota_{2}n_{2}}],\dots,D^{\,{\dagger}}_{\iota_{r}n_{r}}]$ such that for some integers $m_{1},\dots,m_{r},s_{1},\dots,s_{r}\geqslant 0$ we have:
a) $\iota_{j}\geqslant p^{M-1}b^{*}U(s_{j},m_{j})$, where $j=1,\ldots,r$;
b) $r+\sum_{1\leqslant j\leqslant r}s_{j}\geqslant w$ and $\sum_{1\leqslant j\leqslant r}m_{j}\leqslant m$.
Note that for any $w\geqslant 1$, $C_{w}(\mathcal{L}^{\,{\dagger}})\subset\mathcal{L}^{\,{\dagger}}(w)$.
For $m\in\mathbb{Z}_{\geqslant 0}$, let $\mathcal{L}^{\,{\dagger}}[m]_{W_{M}(k)}$ be the $W_{M}(k)$-submodule of $\mathcal{L}^{{\dagger}}_{W_{M}(k)}$ generated by all monomials $[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]:=[\ldots[D^{\,{\dagger}}_{\iota_{1}n_{1}},\ldots],D^{\,{\dagger}}_{\iota_{r}n_{r}}]$ such that $\mathrm{ch}(\iota_{1})+\dots+\mathrm{ch}(\iota_{r})=m$. Denote by $\mathcal{L}^{{\dagger}}[m]$ the $\mathbb{Z}_{p}$-submodule of $\mathcal{L}^{{\dagger}}$ such that $\mathcal{L}^{{\dagger}}[m]\otimes W_{M}(k)=\mathcal{L}^{{\dagger}}[m]_{W_{M}(k)}$.
Definition.
For $\bar{\iota}=(\iota_{1},\dots,\iota_{r})$ set $r(\bar{\iota})=r$, $\mathrm{ch}(\bar{\iota})=\mathrm{ch}(\iota_{1})+\dots+\mathrm{ch}(\iota_{r})$ and $|\bar{\iota}|=\iota_{1}+\dots+\iota_{r}$.
As a result, we have the grading $\mathcal{L}^{\,{\dagger}}=\oplus_{m\geqslant 0}\mathcal{L}^{\,{\dagger}}[m]$ in the category of Lie algebras. For $w\geqslant 1$, set $\mathcal{L}^{\,{\dagger}}(w)[m]=\mathcal{L}^{\,{\dagger}}(w)\cap\mathcal{L}^{\,{\dagger}}[m]$. Then $\mathcal{L}^{\,{\dagger}}(w)=\oplus_{m\geqslant 0}\mathcal{L}^{\,{\dagger}}(w)[m]$.
Remark.
The algebra $\mathcal{L}^{{\dagger}}[0]$ can be naturally identified with the Lie subalgebra of $\mathcal{L}$ from Section 1 via $j^{0}:D^{{\dagger}}_{qp^{M-1}a,n}\mapsto D_{an}$, where $a$ runs over $\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0})$. Then the special choice of $\delta_{0}$ and condition C3 from Section 2.1 imply that this identification is compatible with filtrations, i.e. for any $w\geqslant 1$, $j^{0}(\mathcal{L}^{{\dagger}}(w))=\mathcal{L}(w)$.
Let ¯$\bar{\mathcal{L}}^{\,{\dagger}}=\mathcal{L}^{\,{\dagger}}/\mathcal{L}^{{\dagger}}(p)$. This algebra is provided with the induced central filtration {¯$\{\bar{\mathcal{L}}^{\,{\dagger}}(w)\}_{w\geqslant 1}$ such that ¯$\bar{\mathcal{L}}^{\,{\dagger}}(p)=0$. We also have the induced gradings ¯$\bar{\mathcal{L}}^{\,{\dagger}}=\oplus_{m\geqslant 0}\bar{\mathcal{L}}^{\,{\dagger}}[m]$ and ¯$\bar{\mathcal{L}}^{\,{\dagger}}(w)=\oplus_{m\geqslant 0}\bar{\mathcal{L}}^{\,{\dagger}}(w)[m]$, where ¯$\bar{\mathcal{L}}^{\,{\dagger}}(w)[m]:=\bar{\mathcal{L}}^{\,{\dagger}}(w)\cap\bar{\mathcal{L}}^{\,{\dagger}}[m]$.
Definition.
If $l\in\bar{\mathcal{L}}^{\,{\dagger}}[m]_{W_{M}(k)}$, $l\neq 0$, we set $\mathrm{ch}(l)=m$.
Clearly, for any $m_{1},m_{2}$, [¯$[\bar{\mathcal{L}}^{\,{\dagger}}[m_{1}],\bar{\mathcal{L}}^{\,{\dagger}}[m_{2}]]\subset\bar{\mathcal{L}}^{\,{\dagger}}[m_{1}+m_{2}]$ and we have the induced identification ¯$\bar{\mathcal{L}}^{{\dagger}}[0]=\bar{\mathcal{L}}$ which is compatible with filtrations.
2.4. Lie algebras $\mathcal{N}^{{\dagger}}$ and $\bar{\mathcal{N}}^{{\dagger}}$
Consider $N^{*},r^{*},q$ from Sec. 2.1.
Our approach to the study of the ramification filtration in $\mathcal{G}_{<p,M}$ substantially uses the construction of a totally ramified separable extension $\mathcal{K}^{\prime}$ of $\mathcal{K}$ such that $[\mathcal{K}^{\prime}:\mathcal{K}]=q$ and the Herbrand function $\varphi_{\mathcal{K}^{\prime}/\mathcal{K}}$ has only one edge point $(r^{*},r^{*})$. For a detailed construction of $\mathcal{K}^{\prime}$ cf. e.g. [3], Sect. 1.5. We just need that $\mathcal{K}^{\prime}=k((t_{01}))$, where our fixed local parameter $t_{0}$ of $\mathcal{K}$ equals $t_{01}^{q}E(t_{01}^{b^{*}})^{-1}$ and $E=E(X)$ is the Hasse-Artin exponential.
We use the uniformiser $t_{01}$ to construct (similarly to Sec. 1.1) the corresponding lifts $O^{\prime}_{M}(\mathcal{E}^{\prime})$ of separable extensions $\mathcal{E}^{\prime}/\mathcal{K}^{\prime}$. Notice that for any separable extension $\mathcal{E}^{\prime}$ of $\mathcal{K}^{\prime}$, $O^{\prime}_{M}(\mathcal{E}^{\prime})=W_{M}(\sigma^{M-1}\mathcal{E}^{\prime})[t_{1}]\subset W_{M}(\mathcal{E}^{\prime})$, where $t_{1}=[t_{01}]$ is the Teichmuller representative of $t_{01}$. In particular, $O^{\prime}_{M}(\mathcal{K}^{\prime})=W_{M}(k)((t_{1}))$. There is no natural embedding of $O_{M}(\mathcal{K})$ into $O^{\prime}_{M}(\mathcal{K}^{\prime})$. However, we have the embedding
$\iota:\sigma^{M-1}O_{M}(\mathcal{K})\subset O^{\prime}_{M}(\mathcal{K}^{\prime})$
determined by the “modulo $p^{M}$” lift $t^{p^{M-1}}=t_{1}^{qp^{M-1}}E(t_{1}^{b^{*}})^{-p^{M-1}}\,$ to $O^{\prime}_{M}(\mathcal{K}^{\prime})$ of the “modulo $p$” relation $t_{0}=t_{01}^{q}E(t^{b^{*}}_{01})^{-1}$.
Definition.
Let ${\mathcal{N}}^{\,{\dagger}}$ be the $W_{M}(k)$-submodule in ${\mathcal{L}}^{\,{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ generated by the elements $p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where
a) $b\in\mathfrak{A}^{0}$, $\mathrm{ch}\,b\geqslant\mathrm{ch}(\bar{\iota})$;
b) $-b\geqslant-|\bar{\iota}|+p^{M-1}b^{*}U(\mathrm{ch}(\bar{\iota}),A)\,\,.$
Clearly, $\mathcal{N}^{{\dagger}}$ is a Lie $W_{M}(k)$-subalgebra in $\mathcal{L}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$. We can define similarly $\bar{\mathcal{N}}^{{\dagger}}$ by replacing $\mathcal{L}^{{\dagger}}$ with $\bar{\mathcal{L}}^{{\dagger}}$. The natural projection $\mathcal{L}^{{\dagger}}\to\bar{\mathcal{L}}^{{\dagger}}$ induces the projection $\mathcal{N}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}$.
Definition.
For $j,m\geqslant 0$, let ${\mathcal{N}}^{\,{\dagger}}[j,m]$ be the $W_{M}(k)$-submodule in ${\mathcal{N}}^{\,{\dagger}}$ generated by the elements $p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where
a) $\mathrm{ch}(\bar{\iota})=m$, $\mathrm{ch}\,b=j+m$;
b) $-b\geqslant-|\bar{\iota}|+p^{M-1}b^{*}U(m+j,A)\,.$
Clearly, there is a decomposition $\oplus_{j,m}{\mathcal{N}}^{\,{\dagger}}[j,m]={\mathcal{N}}^{\,{\dagger}}$ and
$[{\mathcal{N}}^{\,{\dagger}}[j_{1},m_{1}],{\mathcal{N}}^{\,{\dagger}}[j_{2},m_{2}]]\subset{\mathcal{N}}^{\,{\dagger}}[j_{1}+j_{2},m_{1}+m_{2}]\,.$
Definition.
For $s\geqslant 0$, ${\mathcal{N}}^{\,{\dagger}}(s):=\oplus_{j+m\geqslant s}{\mathcal{N}}^{\,{\dagger}}[j,m]$ form a decreasing filtration in ${\mathcal{N}}^{\,{\dagger}}$ such that $[{\mathcal{N}}^{{\dagger}}(s_{1}),{\mathcal{N}}^{{\dagger}}(s_{2})]\subset{\mathcal{N}}^{\,{\dagger}}(s_{1}+s_{2})$.
Similarly to Section 1.5 we can introduce for $w\geqslant 1$, a closed submodule $\mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)$ in $\mathcal{L}^{{\dagger}}$ generated by all $p^{m}[D_{\bar{\iota}\bar{n}}]$ with $|\bar{\iota}|\geqslant p^{M-1}b^{*}U(w-1,m)$. Then there are analogs of Proposition 1.6 and Corollary 1.8
$\mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)\cap C_{2}(\mathcal{L}^{{\dagger}}_{W_{M}(k)})\subset\sum_{w_{1}+w_{2}\geqslant w}[\mathcal{L}_{\mathcal{N}}(w_{1}),\mathcal{L}_{\mathcal{N}}(w_{2})],\ \ \mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)\subset\mathcal{L}^{{\dagger}}(w)\,.$
As a result, for any $w$ and $\mathrm{m}^{\prime}=W_{M}(k)[[t_{1}]]$, it holds
$\mathcal{N}^{{\dagger}}(s)\subset\mathcal{L}^{{\dagger}}_{\mathrm{m}^{\prime}}+\mathcal{L}^{{\dagger}}(1+s)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Remark.
Similar definitions can be done in the context of the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$. In particular, we obtain an analog $\bar{\mathcal{N}}^{{\dagger}}$ of $\mathcal{N}^{{\dagger}}$ with corresponding filtration $\bar{\mathcal{N}}^{{\dagger}}(s)$ and grading $\bar{\mathcal{N}}^{{\dagger}}[j,m]$. The projection $\mathcal{N}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}$ allows us to transform above properties to $\bar{\mathcal{N}}^{{\dagger}}$.
In particular, we obtain
Corollary 2.5.
$\bar{\mathcal{N}}^{\,{\dagger}}(p-1)\subset\mathcal{L}^{\,{\dagger}}_{\mathrm{m}^{\prime}}$.
Definition.
Set $\widetilde{\mathcal{N}}^{\,{\dagger}}=\bar{\mathcal{N}}^{\,{\dagger}}/\bar{\mathcal{N}}^{\,{\dagger}}(p)$.
The above grading on $\bar{\mathcal{N}}^{\,{\dagger}}$ induces the grading $\widetilde{\mathcal{N}}^{\,{\dagger}}=\oplus_{j,m}{\widetilde{\mathcal{N}}}^{\,{\dagger}}[j,m]$. Note that $\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]=0$ if $j+m\geqslant p$.
2.5. The action $\Omega$
Suppose $\gamma\in\mathbb{Z}/p^{M}$.
For $b=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ and $l=p^{A}t_{1}^{-b}D^{\,{\dagger}}_{\bar{\iota}\bar{n}}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\,$, set
$\Omega(l):=E(t_{1}^{b^{*}})^{(p^{m}\alpha)p^{M-1}}l\,,$
where $E(X)$ is the Artin-Hasse exponential.
Note that $E(t_{1}^{b^{*}})^{(p^{m}\alpha)p^{M-1}}\equiv\widetilde{\exp}\left((p^{m}\alpha)S_{b^{*}}\right)\mathrm{mod}\,S_{b^{*}}^{p}$ where $S_{b^{*}}=t_{1}^{b^{*}p^{M-1}}+pt_{1}^{b^{*}p^{M-2}}+\ldots+p^{M-1}t_{1}^{b^{*}}$ and $\widetilde{\exp}$ is the truncated exponential. Therefore, in $\widetilde{\mathcal{N}}^{\,{\dagger}}$ it holds $\Omega(l)=\widetilde{\exp}\left((p^{m}\alpha)S_{b^{*}}\right)l$.
Remark.
Note that $S_{b^{*}}^{p}\bar{\mathcal{N}}^{{\dagger}}\subset\bar{\mathcal{N}}^{{\dagger}}(p)$. Therefore, it is essential that we could define the action $\Omega$ only in the context of $\widetilde{\mathcal{N}}^{{\dagger}}=\bar{\mathcal{N}}^{{\dagger}}/\bar{\mathcal{N}}^{{\dagger}}(p)$. However, $\widetilde{\mathcal{N}}^{{\dagger}}$ is still related to the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$, i.e. the natural embedding ¯$\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}\subset\bar{\mathcal{L}}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ still induces embedding ¯$\bar{\mathcal{L}}^{{\dagger}}\to\widetilde{\mathcal{N}}^{{\dagger}}$ after taking the projection from $\bar{\mathcal{N}}^{{\dagger}}$ to $\widetilde{\mathcal{N}}^{{\dagger}}$. (Use that $\bar{\mathcal{N}}^{{\dagger}}(p)\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$.)
By linearity, $\Omega$ can be extended to the whole $\widetilde{\mathcal{N}}^{\,{\dagger}}$.
The map $n\mapsto\Omega_{\gamma}(n):=\Omega^{\gamma}(n)$, $\gamma\in\mathbb{Z}/p^{M}$, is a well-defined action of the additive group $\mathbb{Z}/p^{M}$ on the Lie algebra $\widetilde{\mathcal{N}}^{\,{\dagger}}$. This action is unipotent: for any $n\in\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]$, $\Omega_{\gamma}(n)-n\in\oplus_{i>j}\widetilde{\mathcal{N}}^{\,{\dagger}}[i,m]\,.$
Let $\bar{e}^{\,{\dagger}}\in\bar{\mathcal{N}}^{{\dagger}}$ be such that
(2.1) $\ \ \bar{e}^{\,{\dagger}}\equiv\sum\limits_{\iota\in\mathfrak{A}^{0}(p)}t_{1}^{-\iota}D^{\,{\dagger}}_{\iota 0}\,\mathrm{mod}\,C_{2}(\bar{\mathcal{L}}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})})\,.$
Remark.
Such element is used usually in the nilpotent Artin-Shreier theory, cf. Section 3. In the above expression the elemnt $D^{{\dagger}}_{00}$ is not defined. In this and similar case for the element $e$ from Section 3 we set $D^{{\dagger}}_{00}:=\alpha D^{{\dagger}}_{0}$, where $\alpha=\hat{\alpha}\,\mathrm{mod}\,p^{M}\in W_{M}(k)$ and the trace of $\hat{\alpha}$ over $\mathbb{Q}_{p}$ is equal to 1.
Denote by $\widetilde{e}^{\,{\dagger}}$ the image of $\bar{e}^{\,{\dagger}}$ in $\widetilde{\mathcal{N}}^{\,{\dagger}}$.
A choice of $\bar{e}^{\,{\dagger}}$ allows us to associate to the above defined action $\Omega_{\gamma}$ the “conjugated” action of $A^{\,{\dagger}}_{\gamma}$ on $\bar{\mathcal{L}}^{\,{\dagger}}$ as follows.
Let $\widetilde{\mathcal{N}}^{\,{\dagger}}\langle j\rangle=\oplus_{m}\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]$ and $\widetilde{\mathcal{N}}^{\,{\dagger}}\{s\}:=\oplus_{j\geqslant s}\widetilde{\mathcal{N}}^{\,{\dagger}}\langle j\rangle$.
Proposition 2.6.
For $\gamma\in\mathbb{Z}/p^{M}$, there are unique $\widetilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}$ and $A^{\,{\dagger}}_{\gamma}\in\mathrm{Aut}_{\mathrm{Lie}}\bar{\mathcal{L}}^{\,{\dagger}}$ such that
a) $\sigma\tilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}$ and $\Omega_{\gamma}(\widetilde{e}^{\,{\dagger}})=(\sigma\widetilde{c}_{\gamma})\circ(A^{\,{\dagger}}_{\gamma}\otimes\mathrm{id}_{\mathcal{K}})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}_{\gamma})$;
b) for any $\iota\in\mathfrak{A}^{0}(p)$, $A^{\,{\dagger}}_{\gamma}(D^{\,{\dagger}}_{\iota 0})-D^{\,{\dagger}}_{\iota 0}\in\oplus_{m<\mathrm{ch}(\iota)}\bar{\mathcal{L}}^{\,{\dagger}}[m]_{W_{M}(k)}$.
Proof.
We need the following lemma.
Lemma 2.7.
Suppose $n\in\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ where $j+m\geqslant 1$. Then there are unique $\mathcal{S}(n),\mathcal{R}(n)\in\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ such that:
a) $\mathcal{R}(n)=\sum\limits_{\iota}t^{-\iota}l_{\iota}$ where $\iota$ runs over $\mathfrak{A}^{+}(p)$ with $\mathrm{ch}(\iota)=j+m$ and all $l_{\iota}\in\left(\bar{\mathcal{L}}^{\,{\dagger}}[m]\cap C_{s}(\bar{\mathcal{L}}^{\,{\dagger}})\right)_{W_{M}(k)}$;
b) $n=\mathcal{R}(n)+(\sigma-\mathrm{id})\mathcal{S}(n)$.
Proof of lemma.
It will be sufficient to specify the elements $\mathcal{R}(n)$ and $\mathcal{S}(n)$ for $n=p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where $b\in\mathfrak{A}^{0}$, $\mathrm{ch}(b)=j+m$, $\mathrm{ch}(\bar{\iota})=m$.
It is essential that $\mathrm{ch}(b)\geqslant 1$ and, therefore, $b\neq 0$.
If $b>0$ we set $\mathcal{R}(n)=0$ and $\mathcal{S}(n)=-\sum_{i\geqslant 0}\sigma^{i}(n)$.
If $b=\iota_{0}p^{i_{0}}<0$ with $\iota_{0}\in\mathfrak{A}^{+}(p)$ we set $\mathcal{R}(n)=\sigma^{-i_{0}}n$ and $\mathcal{S}(n)=\sum_{0\leqslant j<i_{0}}\sigma^{-j}n$.
The uniqueness property holds because $\mathrm{ch}(b)\geqslant 1$ implies that $b\neq 0$. Lemma is proved. ∎
Continue the proof of Prop. 2.6.
Use induction on $i\geqslant 1$ to prove the proposition modulo the ideals $\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap C_{i}(\bar{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
— If $i=1$ take $\widetilde{c}_{\gamma}=0$, $A^{\,{\dagger}}_{\gamma}=\mathrm{id}$ and use that $\Omega_{\gamma}(\widetilde{e}^{\,{\dagger}})-\widetilde{e}^{\,{\dagger}}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\,.$
— Suppose $i\geqslant 2$, $\widetilde{c}_{\gamma}\in\widetilde{\mathcal{N}}^{{\dagger}}\{1\}$ and $A_{\gamma}^{\,{\dagger}}\in\mathrm{Aut}_{\mathrm{Lie}}(\bar{\mathcal{L}}^{\,{\dagger}})$ are such that
$H=\Omega_{\gamma}\widetilde{e}^{\,{\dagger}}-(\sigma\widetilde{c}_{\gamma})\circ(A^{\,{\dagger}}_{\gamma}\otimes\mathrm{id}_{\mathcal{K}^{\prime}})\widetilde{\,e}^{\,{\dagger}}\circ(-\widetilde{c}_{\gamma})\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap\,C_{i}(\widetilde{\mathcal{N}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Then $\mathcal{R}(H),\mathcal{S}(H)\in\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap\,C_{i}(\widetilde{\mathcal{L}}^{\,{\dagger}})_{O^{\prime}_{M}(\mathcal{K}^{\prime}}$.
Set $\mathcal{R}(H)=\sum\limits_{\mathrm{ch}(\iota)\geqslant 1+m}t^{-\iota}H_{\iota m}$, where all $H_{\iota m}\in\widetilde{\mathcal{L}}^{\,{\dagger}}[m]_{k}\cap C_{i}(\bar{\mathcal{L}}^{\,{\dagger}})_{W_{M}(k)}$.
Introduce $A^{\,{\dagger}\prime}_{\gamma}\in\mathrm{Aut}_{\mathrm{Lie}}(\bar{\mathcal{L}}^{\,{\dagger}})$ by setting for all involved $\iota$ and $m$, $A_{\gamma}^{\,{\dagger}\prime}(D^{\,{\dagger}}_{\iota 0})=A^{\,{\dagger}}_{\gamma}(D^{\,{\dagger}}_{\iota 0})-\sum_{m}H_{\iota m}$. Set also $\widetilde{c}^{\,\prime}_{\gamma}=\widetilde{c}_{\,\gamma}-{\mathcal{S}}(H)$. Then
$\Omega_{\gamma}\widetilde{e}^{\,{\dagger}}\equiv(\sigma\widetilde{c}^{\,\prime}_{\gamma 1})\circ(A^{\,{\dagger}\prime}_{\gamma}\ \otimes\mathrm{id}_{\mathcal{K}})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}^{\,\prime}_{\gamma})\,\mathrm{mod}\,\widetilde{\mathcal{N}}^{\,{\dagger}}\{1\}\cap C_{i+1}(\widetilde{\mathcal{N}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
The uniqueness follows similarly by induction on $i$ and the uniqueness part of Lemma 2.7.
The proposition is proved. ∎
We have obviously the following properties.
Corollary 2.8.
For any $\gamma,\gamma_{1}\in\mathbb{Z}/p^{M}$,
a) $A^{\,{\dagger}}_{\gamma+\gamma_{1}}=A^{\,{\dagger}}_{\gamma}A^{\,{\dagger}}_{\gamma_{1}}$;
b) $\Omega_{\gamma}(\widetilde{c}_{\gamma_{1}})\circ(A^{\,{\dagger}}_{\gamma_{1}}\otimes\mathrm{id}_{\mathcal{K}})\widetilde{c}_{\gamma}=\widetilde{c}_{\gamma+\gamma_{1}}$;
c) if $l\in\bar{\mathcal{L}}^{\,{\dagger}}[m]$ then $A^{\,{\dagger}}_{\gamma}(l)-l\in\oplus_{m^{\prime}<m}\bar{\mathcal{L}}^{\,{\dagger}}[m^{\prime}]$, e.g. $A^{\,{\dagger}}_{\gamma}|_{\bar{\mathcal{L}}^{\,{\dagger}}[0]}=\mathrm{id}\,.$
2.6. The action $\Omega_{U}$
Let $A^{\,{\dagger}}:=A^{\,{\dagger}}_{\gamma}|_{\gamma=1}$. Then for any $\gamma=n\,\mathrm{mod}\,p^{M}$, $A^{\,{\dagger}}_{\gamma}=A^{\,{\dagger}n}$, in particular, $A^{\,{\dagger}p^{M}}=\mathrm{id}_{\bar{\mathcal{L}}^{\,{\dagger}}}$. By part c) of the above corollary, for all $m\geqslant 0$, there is a differentiation $B^{\,{\dagger}}\in\mathrm{End}_{\mathrm{Lie}}\bar{\mathcal{L}}^{\,{\dagger}}$ such that for all $m\geqslant 0$, $B^{\,{\dagger}}(\bar{\mathcal{L}}^{\,{\dagger}}[m])\subset\oplus_{m^{\prime}<m}\bar{\mathcal{L}}^{\,{\dagger}}[m^{\prime}]$ and for all $\gamma\in\mathbb{Z}/p^{M}$, $A^{\,{\dagger}}_{\gamma}=\widetilde{\exp}(\gamma B^{\,{\dagger}})$.
The derivation $B^{{\dagger}}$ can be recovered via methods from [9]-[10].
Let $U$ be an indeterminate. If $b=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ and $l=p^{A}t_{1}^{-b}[D_{\bar{\iota}\bar{n}}]\in\widetilde{\mathcal{N}}^{\,{\dagger}}\,$ set
$\Omega_{U}l:=\widetilde{\exp}(U\otimes(p^{m}\alpha)S_{b^{*}})l\,\in\mathbb{Z}/p^{M}[U]\otimes\widetilde{\mathcal{N}}^{\,{\dagger}}\,.$
Proceeding similarly to the proof of Proposition 2.6 we obtain
(2.2) $\Omega_{U}(\widetilde{e}^{\,{\dagger}})=\sigma(\widetilde{c}_{U})\circ(A^{\,{\dagger}}_{U}\otimes\mathrm{id})\widetilde{e}^{\,{\dagger}}\circ(-\widetilde{c}_{U})\,,$
where $A^{\,{\dagger}}_{U}=\widetilde{\exp}(UB^{\,{\dagger}})$. As a matter of fact, when obtaining (2.2) we are working inside $\oplus_{m\geqslant 0}U^{j}\widetilde{\mathcal{N}}^{{\dagger}}[j,m]$.
It is easy to see that we have action of the additive group $\mathbb{G}_{a}=\mathrm{Spec}\,\mathbb{Z}/p^{M}[U]$ on $\widetilde{\mathcal{N}}^{\,{\dagger}}$. Similarly to [12] and by taking into account the compatibility of this action with the double grading $\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]$ we obtain the following:
Proposition 2.9.
a) $\widetilde{c}_{U}=\widetilde{c}^{\,(1)}U+\dots+\widetilde{c}^{\,(p-1)}U^{p-1}$, where all $\widetilde{c}^{(j)},\sigma\widetilde{c}^{(j)}\in\widetilde{\mathcal{N}}^{\,{\dagger}}\langle j\rangle$;
b) the cocycle $\widetilde{c}_{U}$ is determined uniquely by its linear part $\widetilde{c}^{\,(1)}$;
c) the action $\Omega_{U}=\sum_{0\leqslant i<p}\Omega^{\,(i)}U^{i}$ (here $\Omega^{(0)}=\mathrm{id}$) is recovered uniquely from its differential $d\,\Omega_{U}:=\Omega^{(1)}U$.
d) $A^{{\dagger}}_{U}=\widetilde{\exp}B^{{\dagger}}_{U}$ and $B^{{\dagger}}_{U}(\bar{\mathcal{L}}^{{\dagger}}[m])\subset\bar{\mathcal{L}}^{{\dagger}}[m-1]$.
Remark.
Prop. 2.9 is directly related to Corollary 2.8 as follows:
1) For any $\gamma\in\mathbb{Z}/p^{M}$, $\widetilde{c}_{U}|_{U=\gamma}=\widetilde{c}_{\gamma}$ and $\Omega_{U}|_{U=\gamma}=\Omega_{\gamma}$.
2) Vice versa, the coefficients $\widetilde{c}^{(j)}$ can be uniquely recovered from the values $\widetilde{c}_{U}|_{U=[\alpha]}$, where $[\alpha]$ are the Teichmuller representatives of $\alpha\in\mathbb{F}_{p}$. (Use that the matrix of Teichmuller representatives $(([i\,\mathrm{mod}\,p]^{j}))_{1\leqslant i,j<p}$, is non-degenerate.)
Corollary 2.10.
For $m\in\mathbb{N}$, $B^{\,{\dagger}}(\widetilde{\mathcal{L}}^{\,{\dagger}}[m])\subset\widetilde{\mathcal{L}}^{\,{\dagger}}[m-1]$.
2.7. Ideals ¯$\bar{\mathcal{L}}^{\,{\dagger}[v_{0}]}$ and ¯$\bar{\mathcal{L}}^{[v_{0}]}$
Recall that ¯$\bar{\mathcal{L}}^{\,{\dagger}}[0]$ is the minimal Lie subalgebra of $\bar{\mathcal{L}}^{\,{\dagger}}$ containing all $D^{\,{\dagger}}_{\iota n}$ with $\mathrm{ch}(\iota)=0$. By Remark from Sect. 2.3, $\mathcal{L}^{{\dagger}}[0]_{W_{M}(k)}$ is generated by
(2.3) $\{D^{{\dagger}}_{qp^{M-1}a,n}\ |\ a\in\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0}),n\in\mathbb{Z}/N_{0}\,\}\,.$
Recall also that $\mathcal{L}^{{\dagger}}[0]$ is naturally identified with the Lie algebra $\mathcal{L}$ via $j^{0}:D^{{\dagger}}_{qp^{M-1}a,n}\mapsto D_{an}$. This identification agrees with filtrations and plays a crucial role below, where we work with criterion from Prop. 3.2 in the algebra $\mathcal{L}^{{\dagger}}[0]$.
[…]
Let $\bar{\mathcal{V}}:=\bar{j^{0}}\bar{\mathcal{V}}^{0}:\bar{\mathcal{L}}^{\,{\dagger}}\longrightarrow\bar{\mathcal{L}}$.
Define the ideal ¯$\bar{\mathcal{L}}^{\,{\dagger}[v_{0}]}$ as the minimal ideal in $\bar{\mathcal{L}}^{\,{\dagger}}$ containing all $B^{\,{\dagger}}(\mathrm{Ker}\,\bar{\mathcal{V}})$. Set ¯$\bar{\mathcal{L}}^{[v_{0}]}=\bar{\mathcal{V}}(\bar{\mathcal{L}}^{\,{\dagger}[v_{0}]})$.
Definition.
Let $\mathfrak{A}^{0}_{1}=\{\iota\in\mathfrak{A}^{0}\ |\ \mathrm{ch}(\iota)=1\}$ and $\mathfrak{A}_{1}^{+}(p)=\mathfrak{A}^{0}_{1}\cap\mathfrak{A}^{+}(p)$.
Proposition 2.11.
¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that for all $\iota\in\mathfrak{A}_{1}^{+}(p)$, $\bar{\mathcal{V}}(D^{{\dagger}}_{\iota 0})\in\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$.
Proof.
Clearly, $\mathrm{Ker}\bar{\mathcal{V}}$ is generated by the elements of ¯$\bar{\mathcal{L}}^{\,{\dagger}}[m]$ with $m\geqslant 1$. Therefore, for $m\geqslant 2$, $B^{\,{\dagger}}(\bar{\mathcal{L}}^{\,{\dagger}}[m])\subset\bar{\mathcal{L}}^{\,{\dagger}}[m-1]\subset\mathrm{Ker}\bar{\mathcal{V}}$. As a result, the elements of $\bar{\mathcal{V}}B^{{\dagger}}(\bar{\mathcal{L}}^{\,{\dagger}}[1])$ generate the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$. If $r(\bar{\iota})>1$ and $\mathrm{ch}(\bar{\iota})=1$ then $l=[D^{{\dagger}}_{\bar{\iota}\bar{n}}]$ can be written in the form $[D^{{\dagger}}_{\iota n},l^{\prime}]$ where $\iota\in\mathfrak{A}_{1}^{+}(p)$ and $l^{\prime}\in\widetilde{\mathcal{L}}^{\,{\dagger}}[0]_{W_{M}(k)}$. This implies that $\bar{\mathcal{V}}B^{\,{\dagger}}(l)\in[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{W_{M}(k)}$.
The proposition is proved. ∎
For each $j\geqslant 1$ set $\bar{\mathcal{N}}^{{\dagger}}\langle j\rangle=\oplus_{m}\bar{\mathcal{N}}^{{\dagger}}[j,m]$, and let $\pi:\bar{\mathcal{N}}\to\widetilde{\mathcal{N}}^{{\dagger}}$ be a natural projection.
Proposition 2.12.
There are unique elements $\bar{c}^{(j)}\in\bar{\mathcal{N}}^{{\dagger}}\langle j\rangle$ such that $\pi(\bar{c}^{(j)})=\widetilde{c}^{(j)}$, cf. Prop. 2.9, and
$\Omega(\bar{e}^{\,{\dagger}})=\sigma(\bar{c})\circ(A^{{\dagger}}\otimes\mathrm{id}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\bar{e}^{\,{\dagger}}\circ(-\bar{c})\,$
where $\bar{c}=\bar{c}^{(1)}+\dots+\bar{c}^{(p-1)}$.
Proof.
It follows easily from the fact that $\mathrm{Ker}\,\pi\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$ and for any $j$ and $m$, the map $\sigma-\mathrm{id}$ is a bijection on $\mathrm{Ker}\pi\cap\bar{\mathcal{N}}^{{\dagger}}[j,m]$. ∎
3. Relation to Galois theory
In [2] we developed a nilpotent analogue of the classical Artin-Schreier theory of cyclic extensions of fields of characteristic $p$. Below we are going to use the covariant analogue of this theory, cf. the discussion in [8], for an explicit description of the ramification subgroup $\mathcal{G}^{(v_{0})}_{<p,M}$ of $\mathcal{G}_{<p,M}=\mathcal{G}/\mathcal{G}^{p^{M}}C_{p}(\mathcal{G})$.
3.1. Groups and Lie algebras of nilpotent class $<p$, [24]
Let $L$ be a finite Lie $\mathbb{Z}/p^{M}$-algebra of nilpotent class $<p$, i.e. $C_{p}(L)=0$. Let $A$ be an enveloping algebra of $L$. There is a natural embedding $L\subset A/J^{p}$ and in terms of this embedding the Campbell-Hausdorff formula appears as
$(l_{1},l_{2})\mapsto l_{1}\circ l_{2}=l_{1}+l_{2}+\frac{1}{2}[l_{1},l_{2}]+\dots,\ \ l_{1},l_{2}\in L\,,$
where $\widetilde{\exp}(l_{1})\cdot\widetilde{\exp}(l_{2})\equiv\widetilde{\exp}(l_{1}\circ l_{2})\,\mathrm{mod}\,J^{p}$. This composition law provides the set $L$ with a group structure and we denote this group by $G(L)$. Clearly, $G(L)$ has period $p^{M}$ and nilpotent class $<p$. Then the correspondence $L\mapsto G(L)$ induces equivalence of the category of finite $p$-groups of period $p^{M}$ and nilpotent class $s<p$ and the category of finite Lie $\mathbb{Z}/p^{M}$-algebras of the same nilpotent class $s$. This equivalence can be naturally extended to similar categories of pro-finite ubjects.
3.2. Epimorphism $\eta^{0}$
Let $L$ be a finite $\mathbb{Z}/p^{M}$-Lie algebra of nilpotent class $<p$. Set $L_{sep}:=L_{O_{M}(\mathcal{K}_{sep})}$. Then the elements of $\mathcal{G}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K})$ and $\sigma$ act on $L_{sep}$ through the second factor, $L_{sep}|_{\sigma=\mathrm{id}}=L$ and $(L_{sep})^{\mathcal{G}}=L_{\mathcal{K}}$. The covariant nilpotent Artin-Schreier theory states that for any $e\in G(L_{\mathcal{K}})$, the set
$\mathcal{F}(e)=\{f\in G(L_{sep})\ |\ \sigma(f)=e\circ f\}$
is not empty and for any fixed $f\in\mathcal{F}(e)$, the map $\tau\mapsto(-f)\circ\tau(f)$ is a continuous group homomorphism $\pi_{f}(e):\mathcal{G}\longrightarrow G(L)$.
Suppose $\mathcal{L}$ is the Lie algebra from Section 1.1 and $e=e(t)\in\mathcal{L}_{\mathcal{K}}$ satisfies $e\equiv\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}\,\mathrm{mod}\,C_{2}(\mathcal{L})_{O_{M}(\mathcal{K})}$. Recall that for $n\in\mathbb{Z}/N_{0}$, $D_{0n}$ is defined as $\sigma^{n}(\alpha_{0})D_{0}$, where $\alpha_{0}=\hat{\alpha}_{0}\,\mathrm{mod}p^{M}$ and $\alpha_{0}\in W(k)$ has absolute trace 1. (We followed this agreement in [1]-[3] to relate the identification $\eta^{0}\,\mathrm{mod}\,C_{2}(\mathcal{G})$, cf. below, to the identification of class field theory.) Then a profinite version of the above Artin-Schreier theory gives the existence of $f\in\mathcal{L}_{sep}$ such that $\sigma f=e\circ f$ and the map $\eta^{0}_{f}:=\pi_{f}(e)\,\mathrm{mod}\,\mathcal{G}^{p^{M}}C_{p}(\mathcal{G})$ induces a group isomorphism $\eta^{0}_{f}:\mathcal{G}_{<p}\simeq G(\mathcal{L})$. We fix a choice of such $f$ and use the simpler notation $\eta^{0}=\eta^{0}_{f}$.
Theorem 3.1 below requires the following assumption:
$\bf A)$ Let $\mathcal{N}$ be a $W_{M}(k)$-Lie subalgera of $\mathcal{L}_{O_{M}(\mathcal{K})}$ generated by all $t^{-a}D_{an}$, $a\in\mathbb{Z}^{0}(p)$. Then $e\in\mathcal{N}$ and $e\equiv\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}\,\mathrm{mod}\,[\mathcal{N},\mathcal{N}]$.
Theorem 3.1.
$\eta^{0}(\mathcal{G}^{(v_{0})})=\mathcal{L}^{[v_{0}]}$.
Remark.
The above quite general result will be explicitly specified in Sec. 4, where we present another proof of the explicit description of $\mathcal{L}^{[v_{0}]}$ in terms of generators $\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p),n\in\mathbb{Z}/N_{0}\}\cup\{D_{0}\}$ of $\mathcal{L}_{W_{M}(k)}$ from [3].
The proof of Theorem 3.1 is given below in sections 3.3-3.5.
3.3. The criterion
Consider the parameters $N^{*},r^{*},q$ from Sec. 2.1.
Consider the field etension $\mathcal{K}^{\prime}/\mathcal{K}$ from Sec. 2.4. Recall $\mathcal{K}^{\prime}=k((t_{01}))$ and the local parameter $t_{01}$ was used to construct the lifts $\ O^{\prime}_{M}(\mathcal{K}^{\prime})$ and $O^{\prime}_{M}(\mathcal{K}_{sep})$. There is a natural embedding $\iota:\sigma^{M-1}O_{M}(\mathcal{K})\to O^{\prime}_{M}(\mathcal{K}^{\prime})$. This allows us to identify the Lie $O^{\prime}_{M}(\mathcal{K}^{\prime})$-algebra $\mathcal{N}^{{\dagger}}[0]$ with subalgebra of $(\mathcal{L}_{\sigma^{M-1}O_{M}(\mathcal{K})})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$. In particular, $\sigma^{M-1}e(t)$ is identified with $\Omega\left(\sigma^{M-1}e(t_{1}^{q})\right)$.
Clearly, we have $\sigma^{M}(f)=\sigma^{M-1}e(t)\circ\sigma^{M-1}f$ and for any $\tau\in\mathcal{G}$,
$\eta^{0}(\tau)=(-\sigma^{M-1}f)\circ\tau(\sigma^{M-1}f)=(-f)\circ\tau(f)\,.$
Choose $f_{1}\in{\mathcal{L}}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ such that $\sigma f_{1}=e(t_{1})\circ f_{1}$ and consider $X\in{\mathcal{L}}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ such that
$\sigma^{M-1}f=X\circ(\sigma^{M-1+N^{*}}f_{1})\,.$
For $Y\in\mathcal{L}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ and an ideal $\mathcal{I}$ in $\mathcal{L}$, define the field of definition of $Y\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ over $\mathcal{K}^{\prime}$ as $\mathcal{K}^{\prime}(Y\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})}):=\mathcal{K}_{sep}^{\mathcal{H}}$, where =$\mathcal{H}=\{g\in\mathcal{G}^{\prime}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K}^{\prime})\ |\ (\mathrm{id}_{\mathcal{L}}\otimes g)Y\equiv Y\,\mathrm{mod}\,\mathcal{I}_{sep}\}$.
For a field extension $\mathcal{E}^{\prime}/\mathcal{K}^{\prime}$ in $\mathcal{K}_{sep}$, define the biggest upper ramification number
$v(\mathcal{E}^{\prime}/\mathcal{K}^{\prime})=\mathrm{max}\{v\ |\ \mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K}^{\prime})^{(v)}\text{acts non-trivially on }\mathcal{E}^{\prime}\}\,.$
The methods from [1, 2, 3] are based on the following criterion.
Set $e_{1}^{(q)}:=e(t_{1}^{q})$. Then $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)=e(t_{1}^{qp^{M-1}}E(t_{1}^{b^{*}})^{-p^{M-1}})$.
Proposition 3.2.
If $X\in\mathcal{L}_{sep}$ is such that
(3.1) $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)=\sigma X\circ\sigma^{M-1}e_{1}^{(q)}\circ(-X)\,$
then $\mathcal{L}^{(v_{0})}$ is the minimal ideal in the set of all ideals $\mathcal{I}$ of $\mathcal{L}$ such that
$v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})})/\mathcal{K}^{\prime})<qv_{0}-b^{*}\,.$
Remark.
The above inequality can be also rewritten in the following way $v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})})/\mathcal{K})<v_{0}\,.$
Suppose $\widetilde{\mathcal{J}}\subset\mathcal{L}$ is a closed ideal and $\pi:\mathcal{L}\longrightarrow L:=\mathcal{L}/\widetilde{\mathcal{J}}$ is a natural projection. Then we can use the appropriate images $e_{L},e_{L}^{(q)}\in L_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ and $f_{L},X_{L}\in L_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ to state the following analog of Prop.3.2 for the ideal $L^{(v_{0})}:=\pi(\mathcal{L}^{(v_{0})})$ of $L$.
Proposition 3.3.
$L^{(v_{0})}$ is the minimal ideal in the set of all ideals $\mathcal{I}$ of $L$ such that $v(\mathcal{K}^{\prime}(X_{L}\,\mathrm{mod}\,\mathcal{I}_{sep})/\mathcal{K}^{\prime})<qv_{0}-b^{*}$.
3.4. Relation between $\mathcal{L}^{(v_{0})}$ and $\mathcal{L}^{[v_{0}]}$
Definition.
For the projection $\mathrm{pr}:\mathcal{L}\to\bar{\mathcal{L}}$, set $\mathcal{L}^{[v_{0}]}:=\mathrm{pr}^{-1}(\bar{\mathcal{L}}^{[v_{0}]})$.
Consider equality (3.1) from the above criterion. On the other hand, according to Proposition 2.12 we have in $\bar{\mathcal{N}}^{\,{\dagger}}$ that
(3.2) $\Omega\left(\sigma^{M-1}\bar{e}^{{\dagger}}\right)=\sigma(\bar{c}_{1})\circ(A^{{\dagger}}\otimes\mathrm{id})\sigma^{M-1}\bar{e}^{{\dagger}}\circ(-\bar{c}_{1})\,.$
Applying $\bar{\mathcal{V}}^{0}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}^{{\dagger}}[0]$ and using the definition of the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ we obtain the following congruence in $\bar{\mathcal{N}}^{{\dagger}}[0]$ with $\bar{x}=\bar{\mathcal{V}}^{0}(\bar{c}_{1})$:
(3.3) $\Omega\left(\sigma^{M-1}\bar{e}_{1}^{(q)}\right)\equiv\sigma(\bar{x})\circ\sigma^{M-1}\bar{e}_{1}^{(q)}\circ(-\bar{x})\,\mathrm{mod}\,\bar{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Since $\mathcal{L}^{{\dagger}}[0](p)=\mathrm{Ker}(\mathcal{L}^{{\dagger}}[0]\to\bar{\mathcal{L}}^{{\dagger}}[0])\subset\mathcal{L}^{[v_{0}]}$ we can take any $x\in\mathcal{N}^{{\dagger}}[0]$ such that $\bar{x}=x\,\mathrm{mod}\,\mathcal{L}^{{\dagger}}[0](p)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ and consider congruence (3.3) in $\mathcal{N}^{{\dagger}}[0]$:
$\Omega\left(\sigma^{M-1}{e}_{1}^{(q)}\right)\equiv\sigma(x)\circ\sigma^{M-1}{e}_{1}^{(q)}\circ(-x)\,\mathrm{mod}\,{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
In particular, the criterion from Prop. 3.2 implies that ${\mathcal{L}}^{(v_{0})}\subset{\mathcal{L}}^{[v_{0}]}$. (Use that $x$ is defined just over a trivial field extension $\mathcal{K}^{\prime}/\mathcal{K}^{\prime}$.)
This means also that we can assume $X\equiv x\,\mathrm{mod}\,\mathcal{L}^{[v_{0}]}_{sep}$.
The following proposition will imply the statement of Theorem 3.1 when $w=p$.
Proposition 3.4.
For any $1\leqslant w<p$, it holds
$a_{w})$  $\mathcal{L}^{(v_{0})}+C_{w}(\mathcal{L})=\mathcal{L}^{[v_{0}]}+C_{w}(\mathcal{L})\,.$
$b_{w})$  $\mathcal{L}(w)\subset\mathcal{L}^{(v_{0})}+C_{w}(\mathcal{L})$.
3.5. The proof of Proposition 3.4
Use induction on $w\geqslant 1$. If $w=1$ there is nothing to prove. So, assume that proposition holds for some $1\leqslant w<p$.
For non-commuting variables $U$ and $V$, let
$\delta^{0}(U,V):=U\circ V-(U+V)\,.$
Note that if the values of $U$ and $V$ are defined modulo $C_{w}(\mathcal{L})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ then the value of $\delta^{0}(U,V)$ is well-defined modulo $C_{w+1}(\mathcal{L})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
Lemma 3.5.
There is a congruence
$\sigma X-X\equiv y\,\mathrm{mod}\,\left([\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L})\right)_{sep}\,,$
where the element $y\in\mathcal{N}^{{\dagger}}[0]$ is such that
$y=\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)-\sigma^{M-1}e_{1}^{(q)}+\delta^{0}\left(\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right),x\right)-\delta^{0}\left(\sigma x,\sigma^{M-1}e_{1}^{(q)}\right)\,.$
Proof of lemma.
Use that $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)\circ X=\sigma X\circ\sigma^{M-1}e_{1}^{(q)}$ and there are the following congruences modulo $\left([\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w_{0}+1}(\mathcal{L})\right)_{sep}$:
$\delta^{0}\left(\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right),X\right)\equiv\delta^{0}\left(\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right),x\right)\,,$
$\delta^{0}(\sigma X,\sigma^{M-1}e_{1}^{(q)})\equiv\delta^{0}\left(\sigma x,\sigma^{M-1}e_{1}^{(q)}\right)\,.$
In order to prove part $b_{w+1})$ we need to verify the following condition:
$(\bullet_{1})$ if $qa\geqslant b^{*}U(w,m)$ then $p^{m}D_{a0}\in(\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L}))_{W_{M}(k)}$.
This will be done by the use of the condition
(3.4) $v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{L}^{(v_{0})}_{sep})/\mathcal{K}^{\prime})<qv_{0}-b^{*}\,.$
Suppose $y=\sum_{\iota}y_{\iota}+y_{+}$, where $y_{+}\in\mathcal{L}_{\mathrm{m}^{\prime}}$, $\iota$ runs over $\mathfrak{A}^{+}(p)$ and the corresponding $y_{\iota}$ appear in the form $y_{\iota}=\sum_{r\geqslant 0}t_{1}^{-\iota p^{r}}l_{br}$ with all $l_{\iota r}\in\mathcal{L}_{W_{M}(k)}$.
Then $X=\sum_{\iota}X_{\iota}+X_{+}$ where $\sigma X_{\iota}-X_{\iota}=y_{\iota}$ and $\sigma X_{+}-X_{+}=y_{+}$. Here $X_{+}\in\mathcal{L}_{\mathrm{m}^{\prime}}$ and does not contribute to (3.4).
The elements $X_{\iota}$ contribute to (3.4) as follows:
if $\iota\in\mathfrak{A}^{+}(p)$ and $\iota/p^{v_{p}(\iota)}\geqslant qv_{0}-b^{*}$ then
$\sum_{r}\sigma^{-r}l_{\iota r}\in(\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L}))_{W_{M}(k)}\,.$
We can follow the contribution to upper ramification in (3.4) for each $\iota\in\mathfrak{A}^{+}(p)$ separately because by Proposiion 2.4 these contributions are independent.
The element $y\in\mathcal{N}^{{\dagger}}[0]\{1\}$ has a linear part
(3.5) $L(y):=\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)-\sigma^{M-1}e_{1}^{(q)}=\sum_{a,s}t_{1}^{-ap^{M-1}q}aS_{b^{*}}^{s}D_{a0}$
and a non-linear part, which is a $W_{M}(k)$-linear combination of various
(3.6) $Q(\iota,\bar{a}_{1},\bar{a}_{2})=p^{A}t_{1}^{-\iota}\left[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]\right]\,,$
where $\iota\in\mathfrak{A}^{0}$, $-\iota\geqslant-qp^{M-1}(|\bar{a}_{1}|+|\bar{a}_{2}|)+p^{M-1-A}b^{*}$. (Recall that $U(1,A)=p^{-A}$, cf. Section 1.)
Note that if the term $Q$ from (3.6) is such that
$|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant(b^{*}/q)U(w,A)$
then $b^{*}/q$ could be replaced by $v_{0}$, cf. Section 2.1, and therefore
$p^{A}\left[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]\right]\in\mathcal{L}_{\mathcal{N}}(w+1)\subset\sum_{w_{1}+w_{2}=w+1}[\mathcal{L}_{\mathcal{N}}(w_{1}),\mathcal{L}_{\mathcal{N}}(w_{2})]\,.$
By inductive assumption $b_{w})$ this term belongs to the $W_{M}(k)$-extension of scalars of
$\sum_{w_{1}+w_{2}=w+1}[\mathcal{L}^{(v_{0})}+C_{w_{1}}(\mathcal{L}),\mathcal{L}^{(v_{0})}+C_{w_{2}}(\mathcal{L})]\subset[\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L})\,.$
This implies that the coefficients for $p^{A}t_{1}^{-\iota}$ in the non-linear part of $y$ with $-\iota\leqslant-p^{M-1}b^{*}U(w,A)$ belong to $[\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L})$ and could be ignored. Indeed, since $-\iota\geqslant-qp^{M-1}(|\bar{a}_{1}|+|\bar{a}_{2}|+p^{M-1}b^{*}p^{M-1-A}$ it holds that $|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant(b^{*}/q)(U(w,m)+b^{*}p^{-m})$. All other terms from the non-linear part will contribute just to smaller upper numbers and also can be ignored. As a result, it remains to care just about the linear part.
If $y_{\iota}=p^{m}t_{1}^{-\iota}l_{\iota}$ is a term from the linear part $L(y)$ such that $\iota>0$ then $\iota\in\mathfrak{A}^{+}(p)$. Therefore, such terms do not appear with $t_{1}^{-\iota p^{r}}$, $r\geqslant 1$. By Prposition 2.4 for different $\iota$, these terms give independent contribution to the upper ramification number in (3.4).
Consider the term $p^{m}t^{-\iota}D_{a0}$ with $-\iota=-aqp^{M-1}+b^{*}p^{M-1}U(w,m)$. If it satisfies the condition $(\bullet_{1})$ then $\iota>0$ and it contributes to the upper number $qp^{i}a-b^{*}(p^{i}U(w,m))$ where $p^{i}U(w,m)\in\mathbb{Z}^{+}(p)$ with $i\geqslant 0$.
Therefore, $qp^{i}a-b^{*}p^{i}U(w,m)>0$ and by Sec.2.1, it is
$\geqslant qp^{i}v_{0}U(w,m)-b^{*}p^{i}U(w,m)=(qv_{0}-b^{*})p^{i}U(w,m)\geqslant qv_{0}-b^{*}\,.$
As a result, $p^{m}D_{a0}\in\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L})$. Part $b_{w+1}$) is proved.
Prove part $a_{w+1}$).
Applying $\bar{\mathcal{V}}^{0}$ to (3.2) we obtain the following congruence in $\bar{\mathcal{N}}^{{\dagger}}[0]$ modulo extension of scalars of the ideal [¯$[\bar{\mathcal{L}}^{{\dagger}[v_{0}]},\bar{\mathcal{L}}^{{\dagger}}]+C_{w+1}(\bar{\mathcal{L}}^{{\dagger}})$:
$\Omega\left(\sigma^{M-1}\bar{e}_{1}^{(q)}\right)\equiv\sigma(\bar{x})\circ(\sigma^{M-1}\bar{e}_{1}^{(q)})\circ(-\bar{x})+\sum_{\iota\in\mathfrak{A}^{+}(p)}t_{1}^{-\iota}\bar{\mathcal{V}}(A^{{\dagger}}-\mathrm{id})D_{\iota 0}\,.$
Comparing it with our criterion
$\Omega\left(\sigma^{M-1}\bar{e}_{1}^{(q)}\right)\equiv\sigma(\bar{X})\circ\sigma^{M-1}\bar{e}_{1}^{(q)}\circ(-\bar{X})\,\mathrm{mod}\,\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{sep}\,.$
we can assume that $\bar{X}=\bar{x}+\bar{Y}$, where $\bar{Y}\in\bar{\mathcal{L}}^{[v_{0}]}_{sep}$ and satisfies the following relation
$\sigma(\bar{Y})-\bar{Y}=\sum_{\iota\in\mathfrak{A}^{+}(p)}t_{1}^{-\iota}\bar{\mathcal{V}}(A^{{\dagger}}-\mathrm{id})(D^{{\dagger}}_{\iota 0})\,.$
Choose the $\iota$-components $\bar{Y}_{\iota}$ such that $\sigma\bar{Y}_{\iota}-\bar{Y}_{\iota}=t_{1}^{-\iota}(A^{{\dagger}}-\mathrm{id})(D_{\iota 0})$ and $\bar{Y}=\sum_{\iota}\bar{Y}_{\iota}$. All $\iota$ give independent contributions to upper ramification numbers. Take $\iota\in\mathfrak{A}_{1}^{+}(p)$, then $\sigma(\bar{Y}_{\iota})-\bar{Y}_{\iota}=t_{1}^{-\iota}B^{{\dagger}}(D_{\iota 0})$.
If $\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})$, cf. Prop. 2.4b), then it gives the upper ramification $q(p^{A}\alpha)-b^{*}$. On the other hand, $\alpha>p^{-A}b^{*}/q$ is equivalent to $\alpha\geqslant p^{-A}v_{0}$ or $qp^{A}\alpha-b^{*}\geqslant qv_{0}-b^{*}$. So, $\bar{\mathcal{V}}B^{{\dagger}}(D_{\iota 0})\in\bar{\mathcal{L}}^{(v_{0})}$.
Therefore, ¯$\bar{\mathcal{L}}^{(v_{0})}+C_{w+1}(\bar{\mathcal{L}})\supset\bar{\mathcal{L}}^{[v_{0}]}+C_{w+1}(\bar{\mathcal{L}}^{[v_{0}]})$.
Theorem 3.1 is completely proved.
3.6. Interpretation in terms of deformations
The above described construction of the ramification ideals $\mathcal{L}^{(v_{0})}$ and ¯$\bar{\mathcal{L}}^{(v_{0})}$ admits the following interpretation if we replace everywhere $t_{1}$ by $t$.
We start, as earlier, with Lie algebra $\mathcal{L}$ and $e=e(t)\in\mathcal{L}_{O_{M}(\mathcal{K})}$ such that (for simplicity) $e=\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}$. Choose $f\in\mathcal{L}_{sep}$ such that $\sigma f=e\circ f$ and fix identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$ via $\tau\mapsto(-f)\circ\tau(f)$, $\tau\in\mathcal{G}_{<p,M}$. As earlier, define the filtration $\mathcal{L}(p)$ and set ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$. Then ¯$\bar{\mathcal{L}}_{W_{M}(k)}$ is generated by $D_{an}$, where $a\in\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0})$. If $\bar{e}\in\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ and $\bar{f}\in\bar{\mathcal{L}}_{sep}$ are the images of $e$ and, resp., $f$ under extensions of scalars of the projection $\mathcal{L}\to\bar{\mathcal{L}}$ then $\tau\mapsto(-\bar{f})\circ\tau(\bar{f})$ determines projection $\bar{\eta}^{0}:\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}})$. Clearly, $\bar{\eta}^{0}$ is induced by $\eta^{0}$.
As earlier, choose $r^{*}$, $N^{*}$, set $q=p^{N^{*}}$, $b^{*}=r^{*}(q-1)$ and introduce the system of coefficients $\{t^{-\iota}\ |\ \iota\in\mathfrak{A}^{0}\}$ from Section 2.2. Consider the quotient ${\mathcal{L}}^{{\dagger}}$ of $\mathcal{L}$ generated by
$\{D_{an}\ |\ a\in[0,(p-1)v_{0}\}\cup\{D_{\iota n}\ |\ \iota\in\mathfrak{A}^{+}(p)\}\subset\{D_{an}\ |\ a\in\mathbb{Z}^{0}(p)\}\,.$
As earlier, define a filtration on $\mathcal{L}^{{\dagger}}$ and set ¯$\bar{\mathcal{L}}^{{\dagger}}=\mathcal{L}^{{\dagger}}/\mathcal{L}^{{\dagger}}(p)$. Note that by Prop. 2.4, all $D_{\iota n}\in\bar{\mathcal{L}}(p)_{W_{M}(k)}$, and therefore, we have the induced projection $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}$.
Take
$e^{{\dagger}}=\sum_{a\in[0,(p-1)v_{0})}t^{-qp^{M-1}a}D_{a,M-1}+\sum_{\iota\in\mathfrak{A}^{+}(p)}t^{-\iota}D_{\iota 0}\,.$
Then the image of $\bar{e}^{{\dagger}}$ in ¯$\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ coincides with $\sigma^{M-1+N^{*}}(\bar{e})$. Therefore, we can choose $f^{{\dagger}}\in\bar{\mathcal{L}}^{{\dagger}}_{sep}$ such that its image in ¯$\bar{\mathcal{L}}_{sep}$ coincides with $\sigma^{M-1+N^{*}}\bar{f}$. This means that we have epimorphism $\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}}^{{\dagger}})$ which gives $\bar{\eta}^{0}$ when composed with projection $G(\bar{\mathcal{V}}):G(\bar{\mathcal{L}}^{{\dagger}})\to G(\bar{\mathcal{L}})$.
Now Theorem 3.1 states that a unipotent $\mathbb{Z}/p^{M}$-action determined by $t^{qp^{M-1}}\mapsto t^{qp^{M-1}}\widetilde{\exp}(-t^{b^{*}})^{p^{M-1}}$ induces via formalism of nilpotent Artin-Schreier theory a deformation $B^{{\dagger}}:\mathrm{Ker}\,\bar{\mathcal{V}}\to\mathrm{Ker}\,\bar{\mathcal{V}}$. This deformation allows us to recover a “substantial” part ¯$\bar{\mathcal{L}}^{(v_{0})}$ of the ramification ideal $\mathcal{L}^{(v_{0})}$. It appears as $\bar{\mathcal{V}}(\mathrm{Ker}\,\bar{\mathcal{V}})$, i.e. as a deformation of the “trivial” part $\mathrm{Ker}\,\bar{\mathcal{V}}\subset\mathcal{L}(p)$ via nilpotent operator $B^{{\dagger}}$.
4. Explicit computations
Now we are going to specify our approach to obtain explicit description of generators of the ramification ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ and to see explicitly how the idea of deformation works in our case.
4.1. Agreements
We need to treat the Campbell-Hausdorff group law via exponentials in the appropriate enveloping algebras $\bar{\mathcal{A}}$ and $\bar{\mathcal{A}}^{{\dagger}}$ of $\bar{\mathcal{L}}$ and, resp., $\bar{\mathcal{L}}^{{\dagger}}$, by involving extension of scalars to $O^{\prime}_{M}(\mathcal{K}^{\prime})$. As a matter of fact, the definition of the action $\Omega$ requires working inside $\widetilde{\mathcal{N}}^{{\dagger}}$ but can be formally treated in $\bar{\mathcal{N}}^{{\dagger}}$, cf. Prop. 2.12. So, it makes sense to work mainly with the Lie algebra ¯$\bar{\mathcal{L}}^{{\dagger}}[0]\simeq\bar{\mathcal{L}}$ and its enveloping algebra $\bar{\mathcal{A}}^{{\dagger}}[0]$.
We are going also to use the advantage of special coefficients from $\mathfrak{A}^{0}$. So, we need just a part of the quotient $(\bar{\mathcal{A}}^{{\dagger}}/\bar{\mathcal{J}}^{{\dagger}p})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ generated by the elements of $\bar{\mathcal{N}}^{{\dagger}}[0]$. (Here $\bar{\mathcal{J}}^{{\dagger}}$ is the augmentation ideal of $\bar{\mathcal{A}}^{{\dagger}}$.) Even more, the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is completely determined by the elements $\bar{\mathcal{V}}^{0}B^{{\dagger}}(D_{\iota 0})$ with $\mathrm{ch}(\iota)=1$. Therefore, it would be reasonable to ignore all $\iota$ with $\mathrm{ch}(\iota)\geqslant 2$, i.e. to work with the part generated by $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}[0]:=\bar{\mathcal{N}}^{{\dagger}}[0,0]\oplus\bar{\mathcal{N}}^{{\dagger}}[1,0]$.
As a result we are going to perform calculations inside the $W_{M}(k)$-module $\bar{\mathcal{A}}^{{\dagger}}_{\leqslant 1}[0]$ generated by all $p^{A}t_{1}^{-\iota}D_{\bar{a}\bar{n}}$, where for $\bar{a}=(a_{1},\dots,a_{r})$ and $\bar{n}=(n_{1},\dots,n_{r})$, $D_{\bar{a}\bar{n}}=D_{a_{1}n_{1}}\ldots D_{a_{r}n_{r}}$, $\iota\in\mathfrak{A}^{0}$, $\mathrm{ch}(\iota)\in\{0,1\}$ and it holds $-\iota\geqslant-qp^{M-1}|\bar{a}|+p^{M-1-A}b^{*}$ (recall that $U(1,A)=p^{-A}$). Since we do not need to follow the effects related to $\mathrm{ch}(\iota)\geqslant 2$, we can agree that $t_{1}^{-\iota}\cdot t_{1}^{-\iota^{\prime}}=0$ if $\iota,\iota^{\prime}\in\mathfrak{A}^{0}_{1}$. We do need to go to $\bar{\mathcal{A}}^{{\dagger}}_{\leqslant 1}[0]$ after applying exponential map to $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}$ but almost immediately we come back to $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}$ and even to $\bar{\mathcal{N}}^{{\dagger}}[1,0]$.
4.2. Choice of $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$
We keep sufficiently general choice of the element $e=e(t)\in{\mathcal{L}}_{O_{M}(\mathcal{K})}$ used for the construction of the identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$. We fix its choice by specifying
$E=\widetilde{\exp}(e)=1+\sum_{r(\bar{a})\geqslant 1}\eta(\bar{a})t^{-(a_{1}+\dots+a_{r})}D_{a_{1}0}\ldots D_{a_{r}0}\,.$
The structural constants $\eta(\bar{a})\in W_{M}(k)$ reflect the fact that $E$ is a “diagonal element” in ${\mathcal{A}}/J^{p}$ or, equivalently, $E\in\widetilde{\exp}(\mathcal{L}_{O_{M}(\mathcal{K})})$ or, equivalently, $\Delta E\equiv E\otimes E$ modulo $p$-th power of augmentation ideal in $\mathcal{A}^{\otimes 2}_{O_{M}(\mathcal{K})}$. Explicitly, this can be stated via the following condition:
$(*_{e})$  if $0\leqslant s_{1}\leqslant s<p$ then
$\eta(a_{1},\dots,a_{s_{1}})\eta(a_{s_{1}+1},\dots,a_{s})=\sum_{\pi\in I_{s_{1}s}}\eta(a_{\pi(1)},\dots,a_{\pi(s)}),$
where $I_{s_{1}s}$ consists of all permutations $\pi$ of order $s$ such that the sequences $\pi^{-1}(1),\dots,\pi^{-1}(s_{1})$ and $\pi^{-1}(s_{1}+1),\dots,\pi^{-1}(s)$ are increasing (i.e. $I_{s_{1}s}$ is the set of all “insertions” of the ordered set $\{1,\dots,s_{1}\}$ into the ordered set $\{s_{1}+1,\dots,s\}$).
Note that in papers [1, 2, 3] we treated on the first place the simplest choice of such constants $\eta(\bar{a})$ related to the case $e=\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}$. In this case the constants are quite simple $\eta(\bar{a})=1/r!$, where $\bar{a}=(a_{1},\dots,a_{r})$. However, our approach works equally well also in the above general situation.
Remark.
There is very interesting possibility to take
$e(t)=D_{00}\circ(t^{-1}D_{10})\circ\ldots\circ(t^{-a}D_{a0})\circ\ldots\,.$
Then $E=\sum_{r(\bar{a})\geqslant 0}\eta(\bar{a})t^{-|\bar{a}|}D_{\bar{a}\bar{0}}$, where $\eta(\bar{a})=0$ if the condition $0\leqslant a_{1}\leqslant a_{2}\leqslant\ldots\leqslant a_{r}$ for $D_{\bar{a}\bar{0}}$ is not satisfied. Note that under this condition $\eta(\bar{a})=1/(r_{1}!r_{2}!\ldots r_{s}!)$ where $\bar{a}=(\bar{a}^{(1)},\ldots,\bar{a}^{(s)})$ and each of the involved vectors $\bar{a}^{(j)}$, $1\leqslant j\leqslant s$, has length $r_{j}$ and equal components. This could be very helpful for studying the ramification breaks of the filtration $\{\mathcal{L}^{(v)}\}_{v\geqslant 0}$.
4.3. Statement of the main result
Definition.
Let $\bar{n}=(n_{1},\dots,n_{s})$ with $s\geqslant 1$. Suppose there is a partition $0=i_{0}<i_{1}<\dots<i_{r}=s$ such that if $i_{j}<u\leqslant i_{j+1}$ then $n_{u}=m_{j+1}$ and $m_{1}>m_{2}>\dots>m_{r}$. Then set
$\eta(\bar{a},\bar{n})=\sigma^{m_{1}}\eta(\bar{a}^{(1)})\dots\sigma^{m_{r}}\eta(\bar{a}^{(r)})\,,$
where $\bar{a}^{(j)}=(a_{i_{j-1}+1},\dots,a_{i_{j}})$. If such a partition does not exist we set $\eta(\bar{a},\bar{n})=0$.
If $s=0$ we set $\eta(\bar{a},\bar{n})=1$. As earlier, set $D_{\bar{a}\bar{n}}=D_{a_{1}n_{1}}\dots D_{a_{s}n_{s}}$.
[…]
For $\gamma\geqslant 0$ and $N\in\mathbb{Z}_{\geqslant 0}$, introduce $\mathcal{F}^{0}_{\gamma,-N}\in{\mathcal{L}}_{W_{M}(k)}$ such that
$\mathcal{F}^{0}_{\gamma,-N}=\sum_{\begin{subarray}{c}\bar{a},\bar{n}\\ \gamma(\bar{a},\bar{n})=\gamma\end{subarray}}a_{1}p^{n_{1}}\eta(\bar{a},\bar{n})[\dots[D_{a_{1}n_{1}},D_{a_{2}n_{2}}],\dots,D_{a_{s}n_{s}}]\,.$
Here:
—  $\bar{a}=(a_{1},\dots,a_{s})$, $\bar{n}\in\mathbb{Z}^{s}$, $n_{1}\geqslant 0$ and all $n_{i}\geqslant-N$;
—  $\gamma=\gamma(\bar{a},\bar{n})=a_{1}p^{n_{1}}+a_{2}p^{n_{2}}+\dots+a_{s}p^{n_{s}}$
Remark.
By definition everywhere $D_{a_{i}n_{i}}:=D_{a_{i},n_{i}\mathrm{mod}N_{0}}$, but when computing $\eta(\bar{a},\bar{n})$ or $\gamma(\bar{a},\bar{n})$ it is essential that $n_{i}\in\mathbb{Z}$.
Note that the non-zero terms in the above expression for $\mathcal{F}_{\gamma,-N}^{0}$ can appear only if $n_{1}\geqslant n_{2}\geqslant\ldots\geqslant n_{s}$ and $\gamma\in A[p-1,N]$, cf. Sect. 2.1.
The result about explicit generators of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ can be stated in the following form.
Let ¯$\bar{\mathcal{F}}^{0}_{\gamma,-N}$ be the image of $\mathcal{F}^{0}_{\gamma,-N}$ in ¯$\bar{\mathcal{L}}_{W_{M}(k)}$.
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}_{1}^{+}(p)$ be the standard presentation from Section 2.2. Note that by definition $\iota\leqslant p^{M-1}(p-1)b^{*}$, $p^{m}\alpha$ and $p^{m}(q-1)\beta/r^{*}$ are integers not both divisible by $p$. We have also that $p^{m}(q-1)\beta_{\iota}/b^{*}=p^{m+M-1-A}$ where $10\leqslant A\leqslant M-$ 1, compare to Prop. 2.4b). So, by setting $\gamma(\iota):=p^{A}\alpha$ and $M(\iota):=m-A$ we obtain the following presentation of $\iota$:
(4.1) $\iota=p^{M-1+M(\iota)}(q\gamma(\iota)-b^{*})\,.$
Remark.
The condition $\iota\leqslant p^{M-1}(p-1)b^{*}$ holds automatically when we work with elements of $\bar{\mathcal{N}}^{{\dagger}}[0]$. Indeed, if $t^{-\iota}D_{\bar{a}\bar{n}}\in\bar{\mathcal{N}}^{{\dagger}}[0]$ then $-\iota\geqslant-qp^{M-1}|\bar{a}|$. Therefore, if $\iota>p^{M-1}(p-1)b^{*}$ then $|\bar{a}|>(p-1)b^{*}/q$. By condition C3 from Section 2.1 it is equivalent to $|\bar{a}|\geqslant(p-1)v_{0}$ and then by Corollary 1.8, $D_{\bar{a}\bar{n}}\in\bar{\mathcal{L}}(p)_{W_{M}(k)}=0$.
Let $m(\iota)$ be the maximal integer such that $\iota p^{m(\iota)}\leqslant p^{M-1}(p-1)b^{*}$. For any $\iota\in\mathfrak{A}^{+}_{1}(p)$, fix a choice of $m_{\iota}\geqslant m(\iota)$.
Theorem 4.1.
¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that for all $\iota\in\mathfrak{A}^{+}_{1}(p)$ given in notation (4.1), ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-(M(\iota)+m_{\iota})}\in\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$.
Remark.
The elements ¯$\bar{\mathcal{F}}_{\gamma,-n}$ appear as linear combinations of Lie monomials $[D_{\bar{a},\bar{n}}]$ which satisfy the condition $\gamma=\gamma(\bar{a},\bar{n})=a_{1}p^{n_{1}}+\ldots+a_{s}p^{n_{s}}$ with $a_{i}\in\mathbb{Z}^{0}(p)$. There is a lot of monomials containing the generators $D_{0n}$ which do not contribute essentially to the condition $\gamma=\gamma(\bar{a},\bar{n})$. It would be natural to switch from the Lie algebra $\mathcal{L}$ to its quotient $\mathcal{L}^{o}=\mathcal{L}/(D_{0})$ which is related just to the generators $\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p)\}$. Then we obtain the description of the corresponding ramification ideal $\mathcal{L}^{o(v_{0})}$ in terms of generators which all contribute (essentially) to the condition $\gamma=\gamma(\bar{a},\bar{n})$. As a matter of fact, this means that we fixed a lift $\hat{\sigma}$ of the Frobenius automorphism to $\mathcal{G}_{<p,M}$ and work with the Galois group of the totally ramified field extension $\mathcal{K}_{<p,M}^{\langle\hat{\sigma}\rangle}$. This field appears very naturally in the traditional local class field theory as well as in its nilpotent version [23].
The proof is given in Sect. 4.4-4.7 below.
4.4. A recurrent relation
Consider the part of relation from Section 3.5 which is adjusted to $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}$
(4.2) $\Omega(\sigma^{M-1}\bar{e}_{1}^{(q)})=\sigma(\bar{x}_{1})\circ\sigma^{M-1}\bar{e}_{1}^{(q)}\circ(-\bar{x}_{1})+\qquad\qquad\qquad$
$\qquad\qquad\qquad\sum_{\iota\in\mathfrak{A}_{1}^{+}(p)}t_{1}^{-\iota}V_{\iota}\,\mathrm{mod}\,[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
where $\bar{x}_{1}\in\bar{\mathcal{N}}^{{\dagger}}[1,0]$ and $V_{\iota}=\bar{\mathcal{V}}^{0}B^{{\dagger}}(D_{\iota 0}^{{\dagger}})$.
Let $E_{0}=\exp(\sigma^{M-1}\bar{e}_{1}^{(q)})$ and $\exp(\Omega(\sigma^{M-1}\bar{e}_{1}^{(q)})=E_{0}+\mathcal{E}_{0}$. In particular, $E_{0}=\sum_{a\in\mathbb{Z}^{0}(p)}\sigma^{M-1}\eta(\bar{a})t_{1}^{-qp^{M-1}|\bar{a}|}D_{\bar{a},M-1}$ and
$\mathcal{E}_{0}=\sum_{r(\bar{a})\geqslant 1}\sigma^{M-1}\eta(\bar{a})|\bar{a}|\sum_{A\geqslant 0}p^{A}t_{1}^{-qp^{M-1}|\bar{a}|+p^{M-1-A}b^{*}}D_{\bar{a},{M-1}}\,.$
Then (4.2) can be rewritten as follows
(4.3) $\sum_{\iota\in\mathfrak{A}_{1}^{+}(p)}t_{1}^{-\iota}V_{\iota}\equiv E_{0}^{-1}\mathcal{E}_{0}+\bar{x}_{1}-E_{0}^{-1}\sigma(\bar{x}_{1})E_{0}\,\mathrm{mod}\,[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Note that this relation is still in ¯$\bar{\mathcal{L}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ due to the following easy lemma:
Lemma 4.2.
If $l\in\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}[0]$, $l^{\prime}\in\bar{\mathcal{N}}^{{\dagger}}[1,0]$ then
$\exp(l+l^{\prime})\cdot\exp(-l)-1\in\bar{\mathcal{N}}^{{\dagger}}[1,0]\,.$
As a result, the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ appears now as the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$ contains the elements $V_{\iota}\in\bar{\mathcal{L}}_{W_{M}(k)}$ from (4.3).
4.5. Solving relation (4.2)
Let $\iota\in\mathfrak{A}^{0}_{1}$. Set
—  $r(\iota):=\max\{r\ |\ \iota p^{r}\in\mathfrak{A}_{1}^{0}\}(=\max\{r\ |\ |\iota p^{r}|\leqslant p^{M-1}(p-1)b^{*}\})$.
—  $\mathfrak{A}_{1}^{\mathrm{prim}}=\mathfrak{A}^{0}_{1}\setminus p\mathfrak{A}^{0}_{1}$. Note that $\mathfrak{A}^{+}_{1}(p)=\{\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}\ |\ \iota>0\}$.
For any $j\geqslant 0$, let $E_{j}=\sigma^{j}E_{0}$ and $\mathcal{E}_{j}=\sigma^{j}\mathcal{E}_{0}$.
Consider the following formal sum in $\bar{\mathcal{A}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$
$\sum_{j\geqslant 0}E_{0}^{-1}E_{1}^{-1}\ldots E_{j}^{-1}\mathcal{E}_{j}E_{j-1}\ldots E_{0}=\sum_{\iota,r}t_{1}^{-\iota p^{r}}\bar{\mathcal{F}}[\iota p^{r}]\,$
where the last sum is taken over all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $r\geqslant 0$.
Proposition 4.3.
For all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $r\geqslant 0$, ¯$\bar{\mathcal{F}}[ip^{r}]\in\bar{\mathcal{L}}_{W_{M}(k)}$.
The proof will be given in Sections 4.6-4.7 below.
For $r\geqslant 0$, let ¯$\bar{\mathcal{F}}[\iota p^{r}]_{0}:=\bar{\mathcal{F}}[\iota p^{r}]-\sigma\bar{\mathcal{F}}[\iota p^{r-1}]\in\bar{\mathcal{L}}_{W_{M}(k)}$. Then
¯$\bar{\mathcal{F}}[\iota p^{r}]=\sum_{0\leqslant j\leqslant r}\sigma^{j}\bar{\mathcal{F}}[\iota p^{r-j}]_{0}$
Proposition 4.4.
a) $\bar{x}_{1}\equiv-\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}\,\mathrm{mod}\,\bar{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}+\bar{\mathcal{N}}^{{\dagger}}(p)$, where the sum is taken over all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $0\leqslant r\leqslant r(\iota)$;
b) if $\iota\in\mathfrak{A}_{1}^{+}(p)$, then
$V_{\iota}\equiv\sigma^{-r(\iota)}\bar{\mathcal{F}}[\iota p^{r(\iota)}]\,\mathrm{mod}\,\left[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}\right]_{W_{M}(k)}$
Proof.
Apply induction on $1\leqslant w<p$ to prove the proposition modulo $C_{w}(\bar{\mathcal{L}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ for all $1\leqslant w\leqslant p$. Clearly, a) holds with $w=1$.
Suppose $1\leqslant w<p$ and a) holds modulo $C_{w}(\bar{\mathcal{L}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
Rewrite (4.2) in the following form
(4.4) $\sum_{\mathrm{ch}\iota=1}t_{1}^{-\iota}V_{\iota}+\sigma{\bar{x}}_{1}^{1}-\bar{x}_{1}^{1}\equiv$
$E_{0}^{-1}\mathcal{E}_{0}-\left(E_{0}^{-1}\sigma\bar{x}_{1}^{1}E_{0}-\sigma\bar{x}^{1}_{1}\right)\,\mathrm{mod}\,[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Now the RHS belongs to $\bar{\mathcal{N}}^{{\dagger}}[1,0]$ modulo extension of scalars of [¯$[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})$. It is easy to see that the RHS appears in the form
$E_{0}^{-1}\mathcal{E}_{0}+\sum_{\iota,r,a}\bar{\mathcal{F}}[qp^{M-1}a+\iota p^{r+1}]_{0}t_{1}^{-(qp^{M-1}a+\iota p^{r+1})}=\sum_{r\leqslant r(\iota)+1,\iota}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\,.$
We can assume that in the last sum $r\leqslant r(\iota)$ due to the following Lemma.
Lemma 4.5.
Let $r\geqslant r(\iota)+1$.
a) If $\iota<0$ then ¯$\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}\in\bar{\mathcal{N}}^{{\dagger}}[0](p)$.
b) If $\iota>0$ then ¯$\bar{\mathcal{F}}[\iota p^{r}]\in\left([\bar{\mathcal{L}}^{{\dagger}[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{W_{M}(k)}$.
Proof of Lemma.
Part a) follows just from definition of $\bar{\mathcal{N}}^{{\dagger}}(p)$. Part b) provided by the inductive assumption about the RHS. ∎
Continue the proof of proposition.
Apply the operators $\mathcal{S}$ and $\mathcal{R}$ from Lemma 2.7 to the RHS of (4.3).
If $\iota<0$ then the expression for $\bar{x}_{1}$ will be provided with
$\mathcal{S}\left(\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=-\sum_{j\geqslant 0}\sigma^{j}\left(\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=-\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]t^{-\iota p^{r}}\,.$
Obviously, the elements from $\bar{\mathcal{N}}^{{\dagger}}[0](p)\subset\bar{\mathcal{L}}_{\mathrm{m}^{\prime}}$ will not contribute to the part $\sum_{\iota\in\mathfrak{A}^{+}(p)}t^{-\iota}V_{\iota}$.
For $\iota>0$, $V_{\iota}t_{1}^{-\iota}$ is congruent modulo ([¯$\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+_{1}}(\bar{\mathcal{L}})\right)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ to
$\mathcal{R}\left(\sum_{r\leqslant r(\iota)}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)\equiv\sum_{r\leqslant r(\iota)}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota}\equiv-\sigma^{-r(\iota)}\bar{\mathcal{F}}[\iota p^{r(\iota)}]\,.$
This proves b).
In this case $\bar{x}_{1}$ will be contributed modulo ¯$\bar{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ with
$\mathcal{S}\left(\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=\sum_{r}\sum_{0\leqslant r_{1}<r}t_{1}^{-\iota p^{r_{1}}}\sigma^{-r+r_{1}}\bar{\mathcal{F}}[\iota p^{r}]_{0}$
$=\sum_{r}t_{1}^{-\iota p^{r}}\sum_{r<r_{1}\leqslant r(\iota)}\sigma^{-r_{1}+r}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}=\sum_{r}t_{1}^{-\iota p^{r}}\sum_{\begin{subarray}{c}r_{1}+r_{2}=r\\ r_{2}<0\end{subarray}}\sigma^{r_{2}}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}$
$\equiv-\sum_{r}t_{1}^{-\iota p^{r}}\sum_{\begin{subarray}{c}r_{1}+r_{2}=r\\ r_{1},r_{2}\geqslant 0\end{subarray}}\sigma^{r_{2}}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}\equiv-\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}$
because
$\sum_{r_{2}+r_{1}=r}\sigma^{r_{2}}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}\equiv\sigma^{r}\left(\sigma^{-r(\iota)}V_{\iota}\right)\,\mathrm{mod}\,\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}}))\right)_{W_{M}(k)}\,.$
4.6. A combinatorial statement
Recall that the values $\eta(\bar{a},\bar{n})$ were defined in Section 4.3. If $e_{(N^{*},0]}:=\sigma^{N^{*}-1}e\circ\sigma^{N^{*}-2}e\circ\dots\circ\sigma e\circ e$ then
$\widetilde{\exp}\,e_{(N^{*},0]}=E_{N^{*}-1}\ldots E_{1}E_{0}=\sum_{\bar{a},\bar{n}}\eta(\bar{a},\bar{n})t_{1}^{-qp^{M-1}\gamma(\bar{a},\bar{n})}D_{\bar{a}\bar{n}}\in\bar{\mathcal{L}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
As earlier, for all $i\geqslant 0$, $E_{i}=\sigma^{i}(E_{0})$ and $\gamma(\bar{a},\bar{n})=a_{1}p^{n_{1}}+\ldots+a_{s}p^{n_{s}}$.
Let $-e_{[0,N^{*})}:=(-e)\circ(-\sigma e)\dots(-\sigma^{N^{*}-1}e)$. Introduce similarly the constants $\eta^{o}(\bar{a},\bar{n})\in W_{M}(k)$ by the following identity
$\widetilde{\exp}(-e_{[0,N^{*})})=\sum\eta^{o}(\bar{a},\bar{n})D_{\bar{a}\bar{n}}\,.$
Set $\eta^{o}(\bar{a}):=\eta^{o}(\bar{a},\bar{0})$. It can be easily seen that if there is a partition from the definition of $\eta$-constants in Sect, 4.3 but such that $m_{1}<m_{2}<\dots<m_{r}$, then
$\eta^{o}(\bar{a},\bar{n})=\sigma^{m_{1}}\left(\eta^{o}(\bar{a}^{(1)})\right)\sigma^{m_{2}}\left(\eta^{o}(\bar{a}^{(2)})\right)\dots\sigma^{m_{r}}\left(\eta^{o}(\bar{a}^{(r)})\right)\,.$
Otherwise, $\eta^{o}(\bar{a},\bar{n})=0$.
If we don’t care what vectors $\bar{a}$ and $\bar{n}$ are we can agree just to write $\eta(1,\dots,s)$ instead of $\eta(\bar{a},\bar{n})$ and use similar agreement for $\eta^{o}$. E.g. the equality $e_{(N^{*},0]}\circ(-e_{[0,N^{*})})=0$ means that
(4.5) $\sum_{0\leqslant s_{1}\leqslant s}\eta(1,\dots,s_{1})\eta^{o}(s_{1}+1,\dots,s)=\delta_{0s}\,.$
For $1\leqslant s_{1}\leqslant s<p$, consider the subset $\Phi_{ss_{1}}$ of permutations $\pi$ of order $s$ such that $\pi(1)=s_{1}$ and for any $1\leqslant l\leqslant s$, the subset $\{\pi(1),\dots,\pi(l)\}\subset[1,s]$ is “connected”, i.e.  for some $n(l)\in\mathbb{N}$,
$\{\pi(1),\dots,\pi(l)\}=\{n(l),n(l)+1,\dots,n(l)+l-1\}.$
By definition, $\Phi_{s0}=\Phi_{s,s+1}=\emptyset$.
Set $B_{s_{1}}(1,\dots,s)=\sum_{\pi\in\Phi_{ss_{1}}}\eta(\pi(1),\dots,\pi(s))$. Then, cf. [12], Sect.4,
Lemma 4.6.
Suppose $0\leqslant s_{1}\leqslant s<p$. Then
a) $B_{s_{1}}(1,\dots,s)+B_{s_{1}+1}(1,\dots,s)=\eta(s_{1},\dots,1)\eta(s_{1}+1,\dots,s)$;
b) $\eta^{o}(1,\dots,s)=(-1)^{s}\eta(s,s-1,\dots,1)$;
c) for indeterminates $X_{1},\dots,X_{s}$,
$\sum_{\begin{subarray}{c}1\leqslant s_{1}\leqslant s\\ \pi\in\Phi_{ss_{1}}\end{subarray}}(-1)^{s_{1}-1}X_{\pi^{-1}(1)}\dots X_{\pi^{-1}(s)}=[\dots[X_{1},X_{2}],\dots,X_{s}].$
4.7. Lie elements ¯$\bar{\mathcal{F}}[\iota]$ and ¯$\bar{\mathcal{F}}[\iota]_{0}$
Agree to write $\bar{n}=(n_{1},\dots,n_{s})\geqslant M$ if all $n_{i}\geqslant M$, and treat similarly $\bar{n}>M$, $\bar{n}\leqslant M$ and $\bar{n}<M$.
For $1\leqslant s_{1}\leqslant s$, let $\gamma^{*}_{[s_{1},s]}(\bar{a},\bar{n})=\sum\limits_{s_{1}\leqslant u\leqslant s}a_{u}p^{n^{*}_{u}}$ where $n_{u}^{*}=0$ if $n_{u}=M(\bar{n}):=\max\{n_{1},\dots,n_{s}\}$ and $n_{u}^{*}=-\infty$ (i.e. $p^{n_{u}^{*}}=0$), otherwise.
Suppose $\iota\in\mathfrak{A}^{0}_{1}$ is given as in (4.1). Then
¯$\bar{\mathcal{F}}[\iota]=\sum_{\bar{a},\bar{n},A}\sum_{1\leqslant s_{1}\leqslant s}p^{A}\eta^{o}(1,\dots,s_{1}-1)\eta(s_{1},\dots,s)\gamma_{[s_{1},s]}^{*}(\bar{a},\bar{n})D_{\bar{a}\bar{n}}$
Here the first sum is taken over all $(\bar{a},\bar{n})$ of lengths $1\leqslant s<p$ such that $\bar{n}\geqslant 0$, $\gamma(\bar{a},\bar{n})=p^{M(\iota)}\gamma(\iota)$ and $M(\bar{n})=M(\iota)+A$.
The following proposition relies elements ¯$\bar{\mathcal{F}}[\iota]$ and the elements ¯$\bar{\mathcal{F}}_{\gamma,-N}$ from Section 4.3.
Proposition 4.7.
If $\iota$ is given in notation (4.1) then
¯$\bar{\mathcal{F}}[\iota p^{n}]=\sigma^{M(\iota)+n}\bar{\mathcal{F}}^{0}_{\gamma(\iota),-(M(\iota)+n)}\,.$
Proof.
We have
$\sigma^{-(M(\iota)+A+n)}\bar{\mathcal{F}}[\iota p^{n}]=\sum_{\begin{subarray}{c}1\leqslant s_{1}\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}p^{A}\eta^{o}(1,\dots,s_{1}-1)\eta(s_{1},\dots,s)\gamma_{[s_{1},s]}^{*}(\bar{a},\bar{n})D_{\bar{a},\bar{n}}\,,$
where the sum is taken for $(\bar{a},\bar{n})$ with $M(\bar{n})=A$, $\bar{n}\geqslant-(M(\iota)+n)$ and $\gamma(\bar{a},\bar{n})=\gamma(\iota)=p^{A}\alpha$. By Lemma 4.2,
$\eta^{o}(1,\dots,s_{1}-1)=(-1)^{s_{1}-1}\eta(s_{1}-1,\dots,1)\,$
and we obtain
$\sum_{\begin{subarray}{c}1\leqslant s_{1}\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}(-1)^{s_{1}-1}(B_{s_{1}-1}(1,\dots,s)+B_{s_{1}}(1,\dots,s))p^{A}\gamma^{*}_{[s_{1},s]}(\bar{a},\bar{n})D_{(\bar{a}\bar{n})}=$
$\sum_{\begin{subarray}{c}1\leqslant s_{1}\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}(-1)^{s_{1}-1}B_{s_{1}}(1,\dots,s)p^{A}(\gamma^{*}_{[s_{1},s]}(\bar{a},\bar{n})-\gamma^{*}_{[s_{1}+1,s]}(\bar{a},\bar{n})){D}_{(\bar{a},\bar{n})}=$
$\sum_{\begin{subarray}{c}1\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}\sum_{1\leqslant s_{1}\leqslant s}(-1)^{s_{1}-1}B_{s_{1}}(1,\dots,s)p^{A}a_{s_{1}}p^{n_{s_{1}}^{*}}{D}_{(\bar{a}\bar{n})}=$
$\sum_{\begin{subarray}{c}1\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}\sum_{\begin{subarray}{c}1\leqslant s_{1}\leqslant s\\ \pi\in\Phi_{ss_{1}}\end{subarray}}(-1)^{s_{1}-1}\eta({\pi(1)},\dots,{\pi(s)})p^{A}a_{s_{1}}p^{n_{s_{1}}^{*}}D_{(\bar{a},\bar{n})}=$
$\sum_{\begin{subarray}{c}1\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}\eta(1,\dots,s)a_{1}p^{A}\sum_{\begin{subarray}{c}1\leqslant s_{1}\leqslant s\\ \pi\in\Phi_{ss_{1}}\end{subarray}}(-1)^{s_{1}-1}D_{a_{\pi^{-1}(1)}n_{\pi^{-1}(1)}}\dots D_{a_{\pi^{-1}(s)}n_{\pi^{-1}(s)}}=$
$\sum_{\begin{subarray}{c}1\leqslant s<p\\ (\bar{a},\bar{n})\end{subarray}}\eta(1,\dots,s)a_{1}p^{n_{1}}[\dots[D_{a_{1}n_{1}},D_{a_{2}n_{2}}],\dots,D_{a_{s}n_{s}}]=\bar{\mathcal{F}}^{0}_{\gamma(iota),-(M(\iota)+n)}\,.$
because $M(\bar{n})=n_{1}$. The proposition is proved. ∎
Corollary 4.8.
All ¯$\bar{\mathcal{F}}[\iota p^{m}]$ and ¯$\bar{\mathcal{F}}[\iota p^{m}]_{0}$ belong to ¯$\bar{\mathcal{L}}_{W_{M}(k)}$.
This finishes the proof of Theorem 4.1.
4.8. Description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ in [3]
Now we use Theorem 4.1 to recover the description of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ obtained in [3].
Theorem 4.9.
Let $\widetilde{N}=N^{*}-1$. Then the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}_{W_{M}(k)}^{[v_{0}]}$ contains all ¯$\bar{\mathcal{F}}^{0}_{\alpha,-\widetilde{N}}$ with $\gamma\geqslant v_{0}$.
Proof.
By Theorem 4.1, ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$ contains all ¯$\bar{\mathcal{F}}[\iota p^{r_{\iota}}]$, where $\iota\in\mathfrak{A}_{1}^{+}(p)$ and $r_{\iota}\geqslant 0$ is such that $\iota p^{r_{\iota}+1}\notin\mathfrak{A}^{0}$.
By (4.1),
$\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})=p^{\widetilde{M}(\iota)+M-1}(q\gamma(\iota)-b^{*})\in\mathfrak{A}_{1}^{+}(p)\,,$
where $\gamma(\iota)=p^{A}\alpha$, $\alpha\in A[p-1,m]$ and $\widetilde{M}(\iota)=m-A$. Note that $\iota\in\mathfrak{A}^{+}_{1}(p)$ means also that $\iota\leqslant p^{M-1}(p-1)b^{*}$ and we have the following two cases (cf. Prop. 4.10 below): either $v_{p}(\gamma(\iota)+\widetilde{M}\iota)=0$ or $\widetilde{M}(\iota)+M-1=0$ (what is equivalent to $A=M-1$ amd $m=0$).
Then by Prop. 4.7, ¯$\bar{\mathcal{F}}[\iota p^{r_{\iota}}]$ can be replaced by ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-\widetilde{N}}$ because $\widetilde{N}=N^{*}-1=\widetilde{M}(\iota)+r_{\iota}$ for some $r_{\iota}\geqslant r(\iota)$, where as earlier
$r(\iota)=\max\{r\geqslant 0\ |\ \iota p^{r(\iota)}\in\mathfrak{A}^{0}\}\,.$
As a result, it remains to verify the following proposition. ∎
Proposition 4.10.
Suppose $\gamma\geqslant v_{0}$ is such that for any $\iota\in\mathfrak{A}^{+}_{1}(p)$, $\gamma\neq\gamma(\iota)$ and $v_{p}(\gamma)\geqslant-\widetilde{N}$. Then ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}\in\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$.
Proof.
As it was mentioned above, we have the two following cases:
1) $p^{M-1-v_{p}(\gamma)}(q\gamma-b^{*})>p^{M-1}(p-1)b^{*}$;
2) $v_{p}(\gamma)\geqslant 0$ and $q\gamma-b^{*}>p^{M-1}(p-1)b^{*}$.
The expression for ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}$ is a linear combination of the terms
$p^{n_{1}}[\dots[D_{a_{1}n_{1}},D_{a_{2}n_{2}}],\dots,D_{a_{r}n_{r}}]\,,$
where $M>n_{1}\geqslant n_{2}\geqslant\ldots\geqslant n_{r}\geqslant-\widetilde{N}$ and $p^{n_{1}}a_{1}+\dots+p^{n_{r}}a_{r}=\gamma$.
We can assume that all $a_{i}<(p-1)v_{0}$, because otherwise ${\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}\in\mathcal{L}(p)_{W_{M}(k)}$ and ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}=0$.
Consider the first case.
From definitions it follows that ¯$\bar{\mathcal{F}}^{0}_{\gamma,-\widetilde{N}}$ is a linear combination of the terms of the form
[[¯$[\dots[\bar{\mathcal{F}}^{0}_{\gamma^{\prime},-(\widetilde{N}-1)},D_{a_{1}^{\prime},-\widetilde{N}}],\dots,D_{a_{r}^{\prime},-\widetilde{N}}]$
where $\gamma=\gamma^{\prime}+(a_{1}^{\prime}+\dots+a_{r}^{\prime})/p^{\widetilde{N}}$, $r\geqslant 0$, all $a^{\prime}_{i}\in\mathbb{Z}^{0}(p)$ and $\gamma^{\prime}\in A[p-1,\widetilde{N}-1]$.
If $a_{1}^{\prime}+\dots+a_{r}^{\prime}\geqslant(p-1)b^{*}/q$, i.e. it is $\geqslant(p-1)v_{0}$, then the whole term belongs to ¯$\bar{\mathcal{L}}(p)_{W_{M}(k)}=0$.
Otherwise, the inequality
$b^{*}(p-1)/p^{\widetilde{N}}<q\gamma-b^{*}=q\gamma^{\prime}-b^{*}+q(a_{1}^{\prime}+\dots+a^{\prime}_{r})/p^{\widetilde{N}}$
implies that $q\gamma^{\prime}-b^{*}>0$ and, therefore, $\gamma^{\prime}\geqslant v_{0}$. If $\gamma^{\prime}=\gamma(\iota^{\prime})$ for a suitable $\iota^{\prime}\in\mathfrak{A}^{+}_{1}(p)$ then ¯$\bar{\mathcal{F}}^{0}_{\gamma^{\prime},-(\widetilde{N}-1)}\in\bar{\mathcal{L}}^{[v_{0}]}$ and we are done. Otherwise, we should repeat the procedure with $\gamma^{\prime}$ instead of $\gamma$.
As a result we either prove our proposition or will come to the second case.
Here we proceed similarly using the above calculations with $\widetilde{N}=0$:
$\gamma=\gamma^{\prime}+(a_{1}^{\prime}+\ldots+a^{\prime}_{r})$, $r\geqslant 0$ and $v_{p}(\gamma^{\prime})\geqslant 1$. Then
$q\gamma-b^{*}=q\gamma^{\prime}-b^{*}+q(a_{1}^{\prime}+\ldots+a^{\prime}_{r})>p^{M-1}(p-1)b^{*}\,.$
Again, if $a_{1}^{\prime}+\ldots+a^{\prime}_{r}\geqslant(p-1)v_{0}$ we are in ¯$\bar{\mathcal{L}}(p)_{W_{M}(k)}$. Otherwise, $q\gamma^{\prime}-b^{*}>0$ implies $\gamma^{\prime}\geqslant v_{0}$ and we have alternative: $\gamma^{\prime}=\gamma(\iota^{\prime})$ with $\iota^{\prime}\in\mathfrak{A}_{1}^{+}(p)$ or repeat the procedure. ∎